### v1

# Urban Region Pre-training and Prompting: A Graph-based Approach

Abstract.
Urban region representation is crucial for various urban downstream tasks. However, despite the proliferation of methods and their success, acquiring general urban region knowledge and adapting to different tasks remains challenging. Previous work often neglects the spatial structures and functional layouts between entities, limiting their ability to capture transferable knowledge across regions. Further, these methods struggle to adapt effectively to specific downstream tasks, as they do not adequately address the unique features and relationships required for different downstream tasks. In this paper, we propose a Graph-based Urban Region Pre-training and Prompting framework (GURPP) for region representation learning. Specifically, we first construct an urban region graph that integrates detailed spatial entity data for more effective urban region representation. Then, we develop a subgraph-centric urban region pre-training model to capture the heterogeneous and transferable patterns of interactions among entities. To further enhance the adaptability of these embeddings to different tasks, we design two graph-based prompting methods to incorporate explicit/hidden task knowledge. Extensive experiments on various urban region prediction tasks and different cities demonstrate the superior performance of our GURPP framework. The implementation is available at this repository: https://anonymous.4open.science/r/GURPP.
Urban Computing, Urban Region Representation Learning
copyright: acmlicensed
1. Introduction
An urban region is a geographical area comprising diverse spatial entities (e.g.,  shopping centers, restaurants, and roads) and their interactions (e.g.,  nearby and located in). Recently, urban region representation learning, which aims to extract vector embeddings from these entities, has become a significant research focus (Zhang et al., 2024). Effective urban region representations can be applied to various tasks, such as region popularity prediction (Li et al., 2024; Fu et al., 2019; Zhou et al., 2023a; Wu et al., 2022), house price prediction (Wang and Li, 2017), crime prediction (Wang and Li, 2017; Zhou et al., 2023a; Wu et al., 2022; Zhang et al., 2021), and population density inference (Li et al., 2023), enabling urban planners, policymakers, and businesses to enhance decision-making processes. However, the complexity of urban environments and the diversity of applications pose significant challenges.
Figure 1. Example of urban generality and task specificity.
Balancing urban generality and task specificity is crucial in urban region representation. (i) Urban regions often appear similar in their spatial structures and functional layouts, because the interactions among entities within these regions follow similar patterns. These shared patterns serve as generic knowledge across urban areas, facilitating the development of generalizable region representation models. As illustrated in Figure 1, Region A and Region B showcase typical urban regions. In these examples, spatial structures are demonstrated by the proximity of residences to parks and shops, which enhances community convenience and accessibility. Functional layouts are characterized by the clustering of similar entities within areas, such as shops in commercial areas and schools in teaching areas, and the close proximity of these distinct areas to each other. Therefore, an effective urban region representation model needs to capture the heterogeneous and generic patterns of interactions among entities within regions. (ii) Despite the broad similarities, each urban region and downstream task possesses unique details. Specific tasks may require attention to different features within urban data. For example, in traffic crash prediction, the density of roads and the number of residents living nearby, might be critical features, as they directly influence the probability of accidents occurring. On the other hand, crime prediction models might focus on the density and type of POIs within the region, such as bars or late-night convenience stores. As consequence, balance between generality and specificity raises a crucial question: Can we design an urban region representation framework that effectively integrates generalizable knowledge while accommodating the specific needs of distinct tasks?
Recently, the pre-training and prompting paradigm that learns general representations that can be fine-tuned or prompted to adapt to specific tasks has shown remarkable success in the field of natural language processing (Liu et al., 2023b; Brown et al., 2020; Roberts et al., 2020; Gao et al., 2021) and computer vision (Radford et al., 2021; Zhou et al., 2022b, a). Inspired by these advancements, we hope to design a pre-training and prompting framework for urban region representation learning. However, building such a framework for urban representation is never easy. The challenges are as follows.
(1) Capturing the heterogeneous and generic patterns of interactions among spatial entities within regions is hard. Most previous methods follow a process where they first predefine aggregations and statistics about urban entities and relationships, and then learn from these predefined features (Han et al., 2020; Yang et al., 2022; Zhang et al., 2023; Li et al., 2023). While this approach has its merits, it often fails to capture the spatial structures and functional layouts between entities within regions which is generalizable. Other previous methods consider urban entities in isolation, but they focus on only a few types, such as POIs and roads and a few relations between them. This narrow focus limits their ability to capture the heterogeneous and generic patterns of interactions among entities, resulting in embeddings that are less robust and less effective (Fu et al., 2019).
(2) Tuning pre-trained region embeddings to capture the tasks’ focus on different features or relationships within the original urban data poses a challenge. The majority of previous urban region representation methods use simple regression models which merely projects the pre-train embeddings without capturing the specific entities and relationships relevant to individual tasks (Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021). In contrast, prompting offers a more flexible approach by allowing models to focus on specific aspects of the data as needed (Liu et al., 2023b; Yoo et al., 2023; Sun et al., 2023), effectively highlighting small but crucial differences between urban downstream tasks. Despite this potential, existing prompting methods (Yan et al., 2024; Zhou et al., 2023a) have limitations. These methods use simple prompts that do not consider detailed and heterogeneous interactions among various urban entities. Additionally, many existing prompting techniques simply concatenate prompts onto the original embeddings, lacking the flexibility to selectively filter out entities and relationships that are specifically useful for downstream tasks. This limitation makes it challenging to tailor the generalizable embedding precisely to the unique characteristics of different urban downstream tasks.
Figure 2. Main idea of urban pre-training and prompting
In light of the challenges discussed, we propose a novel framework named GURPP (Graph-based Urban Region Pre-training and Prompting Framework) as shown in Figure 2. Specifically:
To address the first challenge, we develop the Subgraph-centric Urban Region Pre-training (GURP) model. To capture the spatial structures and functional layouts between entities, we construct an urban region graph that integrates urban entities and their relationships. Unlike previous methods where each region is treated as a single point, our approach views each region as a subgraph. So the GURP model then extract region subgraphs from urban region graph with graph patterns and learns region embeddings from these subgraphs through a heterogeneous graph learning model. To further enhance the learning process, we design a learning pipeline that incorporates heterogeneous graph learning model and multi-view learning. It enables the model to learn both structure information within subgraphs and the relationships between subgraphs.
To address the second challenge, we develop a knowledge-enriched prompting learning model. This model includes two approaches: a manually-designed prompt method and a task-learnable prompt learning method. The manually-designed prompt adjusts subgraph structures to highlight task-specific relationships and entities using predefined weights. The task-learnable prompt learns from task data to generate prompt graphs that capture implicit task-specific knowledge. By incorporating both predefined and learned task-specific prompts, the model tailors the general embeddings to better suit the unique characteristics and requirements of various urban computing tasks.
Contributions:
  • (Urban region graph). We introduce urban region graph, which integrates detailed spatial and entity data for more effective urban region representation.
  • (Pre-training model). We propose a Subgraph-centric Urban Region Pre-training (GURP) model for learning embeddings from urban region graphs, capturing generalizable patterns of interactions among entities within regions.
  • (Prompting model). We design a novel knowledge-enriched prompting learning model with two graph-based prompting methods: a manually-defined prompt to incorporate explicit task knowledge and a task-learnable prompt to discover hidden task-relevant patterns.
  • (Extensive experiments). We demonstrate the effectiveness of our GURPP framework through extensive experiments on three urban region prediction tasks across two cities, achieving significant performance improvements with an average increase of $17.28\%$ on MAE and $22.26\%$ on RMSE.
2. Preliminaries
Urban region. A city is divided into a set $\mathcal{R}$ of non-overlapping regions based on defined criteria, such as administrative boundaries (Zhang et al., 2021; Wu et al., 2022; Liu et al., 2023a; Ning et al., 2024) and grid partitioning (Jin et al., 2024; Yuan et al., 2024; Liu et al., 2019; Feng et al., 2019). Each region $r\in\mathcal{R}$ consists of multiple spatial entities, including points of interest (POIs), roads, and junctions.
Urban region representation learning. The urban region representation learning problem involves learning a mapping function $\zeta:\mathcal{R}\rightarrow\mathbb{R}^{d}$, which generates a low-dimensional embedding $\vec{h}\in\mathbb{R}^{d}$ for each region $r\in\mathcal{R}$. This mapping considers both the spatial structures and functional layouts of the spatial entities within $r$, with $d$ being the size of the embedding.
Once the mapping function $\zeta(\cdot)$ has been learned, it can be applied to various downstream prediction tasks, such as crime and check-in predictions. Let the task dataset be $\{(r_{i},y_{i})\}_{i=1}^{|\mathcal{R}|}$, where $y_{i}\in\mathbb{R}$ is a numerical indicator for region $r_{i}$. The downstream prediction tasks aim to predict $y_{i}$ based on the region embedding $\zeta(r_{i})$ using a simple regression model like a Ridge Regression (Hoerl and Kennard, 1970).
Figure 3. The Overview of our GURPP. (1) Create urban region graph that integrates specific urban entities. (2) Extract subgraph of each region according to the graph pattern. (3) Use multi-view learning to learn the subgraph representations with GURP model. (4) Adapt the region embeddings with manually-designed prompt method and task-learnable prompt learning method.
3. Overview
As shown in Figure 3, we propose a graph-based framework to pre-train and prompt urban region representations. Our approach begins with the creation of an urban region graph that integrates specific urban entities, providing a comprehensive structure for representing spatial and functional relationships within urban areas (discussed in Section 4.1). From this urban region graph, we extract subgraphs corresponding to each region, using graph patterns that characterize the region’s structure and interactions. These subgraph representations are then refined through multi-view learning using the GURP model, enabling the capture of generalizable information that can serve as the foundation for downstream tasks (explained in Section 4.2). To tailor these representations for specific tasks, we employ two prompt-tuning methods: a manually-designed prompt method and a task-learnable prompt learning method, each designed to incorporate task-specific knowledge into the pre-trained region embeddings (elaborated in Sections 5.1 and 5.2, respectively).
4. Pre-train General Representation
In this section, we introduce the processes of constructing the urban region graph and pre-training subgraph-centric region embeddings using a multi-view learning approach.
4.1. Urban Region Graph Construction
An urban region graph is constructed to model the spatial structures and functional layouts of urban areas. The urban region graph, denoted as $G=(V,E,T_{V},T_{E})$, is a heterogeneous graph containing multiple types of nodes and relationships. Here, $v\in V$ and $e\in E$ represent a node and an edge, respectively, and $\phi(v)\in T_{V}$ $(\psi(e)\in T_{E}$) is a type function that assign a type to a specific node (edge).
The urban region graph follows a schema that defines the relationships among various types of nodes and edges (see Figure 4). Without loss of generality, we consider three types of spatial entities: POIs, roads, and junctions. Each spatial entity is categorized: for instance, POI categories include shops and parks, road categories encompass highways and residential streets, and junction categories cover features like roundabouts.
Figure 4. Schema of urban region graph.
We construct the urban region graph from POI datasets and online maps. Since the urban region graph contains heterogeneous nodes and edges, we adopt the knowledge graph embedding technique TransR (Lin et al., 2015) to initialize the representations of all nodes and edges. This aligns the heterogeneous information into a unified vector space.
4.2. Subgraph-centric Region Embedding
We introduce a subgraph-centric region embedding model to encode the spatial structures and functional layouts of a region $r$.
4.2.1. Extract region-induced subgraph.
The region-induced subgraph is a subgraph of $G$ that includes the region node $r$ and matches the graph pattern $\mathcal{P}=(P_{V},P_{E},\theta)$. Here, $P_{V}\subseteq T_{V}$ and $P_{E}\subseteq T_{E}$ are sets of node and edge types, respectively, which are used to filter entities and relationships specifically useful for region embeddings. The set $\theta\subseteq T_{V}$ indicates the node types that serve as termination points for subgraph extraction. As a result, the schema of the region-induced subgraph is a subset of the schema of the urban region graph. During the pre-training phase, we select all node and edge types and set $P_{V}=T_{V},P_{E}=T_{E}$, i.e., $\mathcal{P}=(T_{V},T_{E},\theta)$ with the aim of comprehensively characterising the rich information of a region and thereby learning general knowledge.
The region-induced subgraph $g_{r}=(V_{r},E_{r},P_{V},P_{E})$ includes nodes $V_{r}$ and edges $E_{r}$ that can be reached from $r$ and whose types match $P_{V}$ and $P_{E}$. We extract $g_{r}$ recursively, starting with $V_{r}=\{r\}$ and $E_{r}=\varnothing$ and repeatedly expanding $V_{r}$ and $E_{r}$ until the termination points are reached. Specifically, we use a breadth-first search (BFS) starting from $r$ to expand the subgraph. During each iteration of BFS, we explore edges $e(u,v)$ where one endpoint $u$ is in $V_{r}$ and the other endpoint $v$ is not yet in $V_{r}$. We then check whether the type of $e$ is in $P_{E}$ and whether the type of the unseen node $v$ is not in the termination type set $\theta$. If $e$ passes the checks, we add it to the set $\Delta E_{r}$. Using $\Delta E_{r}$, we construct the node set $\Delta V_{r}$, which contains the endpoints in $\Delta E_{r}$ that are not already in $V_{r}$. Formally, the sets of incremental edges (i.e.,  $\Delta E_{r}$) and nodes (i.e.,  $\Delta V_{r}$) are as follows: $\Delta E_{r}=\{e(u,v)~{}|~{}u\in V_{r}\land\phi(v)\notin\theta\land e\in E\land\psi(e)\in P_{E}\}$ and $\Delta V_{r}=\{v|~{}v\notin V_{r}\land\exists u\in V_{r},e(u,v)\in\Delta E_{r}\}$. We then update the region-induced subgraph by setting $V_{r}=V_{r}\cup\Delta V_{r}$ and $E_{r}=E_{r}\cup\Delta E_{r}$. The expansion terminates if $\Delta E_{r}$ and $\Delta V_{r}$ are empty.
4.2.2. Region subgraph encoding.
In Section 4.1, each node in the subgraph $g_{r}$ has been assigned a global initial embedding using TransR. Here, we encode each subgraph $g_{r}$ into a vector $\vec{h}_{r}$ to further incorporate the local features of region $r$. Specifically, we use a heterogeneous graph encoder $\textsf{gEnc}(\cdot)$ to extract the heterogeneous and generic patterns within $g_{r}$, ultimately outputting the contextualized embedding of the subgraph.
Our encoder $\textsf{gEnc}(\cdot)$ chooses the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) as the subgraph node encoder due to HGT’s capability of capturing node and edge-types semantics. By using HGT, $\textsf{gEnc}(\cdot)$ characterizes the heterogeneity of input graph by maintaining a set of node and edge-type dependent parameters, and the output embedding of each node $\vec{v}^{HGT}$ that integrates the global initial embeddings generated by TransR, and node and edge-type semantics of the subgraph $g_{r}$. This trait aligns with our need to capture the semantic features of heterogeneous entities and relationships within each region-induced subgraph.
After obtaining the representations of each node in the subgraph, we further obtain the graph-level representation of $g_{r}$ through two steps of aggregation. For each subgraph $g_{r}$, we first use $\text{SUM}(\cdot)$ function to aggregate the representations of nodes of the same type to integrate the impact of the number of entities on the representation, obtaining the node-type embedding $\vec{\tau}_{i}$:
$\vec{\tau}_{i}=\sum_{\phi(v)=\tau_{i}}\vec{v}^{HGT},\quad\tau_{i}\in T_{V},\forall v\in V_{r}.$
Then we concatenate all representations of node types, and encode them with a Linear layer to capture the interaction between various types of nodes:
$\vec{h}_{r}=\text{Linear}\left(\text{CONCAT}_{|T_{V}|}(\vec{\tau}_{i})\right).$
The output $\vec{h}_{r}$ of $\textsf{gEnc}(\cdot)$ is the final embedding of region-induced subgraph $g_{r}$. Once the graph encoder is pretrained, any graph that conforms to the pattern $\mathcal{P}$ can be encoded by $\textsf{gEnc}(\cdot)$ to generate the corresponding representations, demonstrating high flexibility.
4.3. Multi-view Representation Learning
We learn the region representation with multiple views to characterize general region features by using the spatial proximity, region satellite imagery and region-wise traffic information.
4.3.1. Spatial proximity.
The First Law of Geography (Tobler, 1970) emphasizes the concept of spatial proximity, stating that “everything is related to everything else, but near things are more related than distant things”, which means two regions with spatial proximity should exhibit greater similarity in their representation.
We utilize a triplet network (Hoffer and Ailon, 2015) to capture region similarity based on spatial proximity. For a given subgraph $g_{r}$ of region $r$, which serves as the anchor sample, we construct its positive and negative samples based on spatial proximity. If region $r^{\prime}$ is adjacent to $r$ (i.e.,  there is a “NearBy” type of edge connecting the corresponding nodes), we take the subgraph of $r^{\prime}$ as the positive sample, denoted by $g_{r}^{+}$. Otherwise, we take the subgraph as the negative sample, denoted by $g_{r}^{-}$. We input $g_{r}$, $g_{r}^{+}$, and $g_{r}^{-}$ to the graph encoder $\textsf{gEnc}(\cdot)$, obtaining the vector representations $\vec{h}_{r}$, $\vec{h}_{r}^{+}$, and $\vec{h}_{r}^{-}\in\mathbb{R}^{d}$, respectively. Then we use triplet loss to minimize the distance between adjacent regions and maximize it between distant regions, with the loss function calculated:
$\mathcal{L}_{sp}=\sum_{r\in\mathcal{R}}\max\{||\vec{h}_{r}-\vec{h}_{r}^{+}||_{2}-||\vec{h}_{r}-\vec{h}_{r}^{-}||_{2}+\delta,0\}.$
where $\delta$ represents the margin parameter of triplet loss.
4.3.2. Region imagery.
Urban imagery encompasses various types of image information that describe the visual characteristics of a city, such as satellite images, street view images, and more. In this paper, we use satellite images as the source of urban imagery. For a given region $r$, its imagery can be represented as the set $\{a_{1},a_{2},\cdots,a_{m}\}$. We employ a pre-trained visual model to extract visual features of the region $\vec{h}_{r}^{img}$ and fuse these features with the region’s representation with the contrastive loss $\mathcal{L}_{img}$ (see Appendix A.2 for more details).
4.3.3. Human mobility.
Human mobility describes the dynamics of urban systems and is typically represented by trajectory data. For a given region $r$, the outflow feature $\mathbf{s}_{r}$ and inflow feature $\mathbf{t}_{r}$ are defined by calculating the number of outgoing and incoming trips within each time period. This results in the outflow and inflow feature sets $\{\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{N}\}$ and $\{\mathbf{t}_{1},\mathbf{t}_{2},\cdots,\mathbf{t}_{N}\}$, respectively, where $\mathbf{s}_{i},\mathbf{t}_{i}\in\mathbb{R}^{l}$, and $l$ is the number of intervals (e.g., 24). We then encode the inflow and outflow features using an encoder (such as a Multi-Layer Perceptron) to obtain $d$-dimensional embeddings:
$\vec{h}^{\text{src}}_{r}=\textsf{Encoder}(\mathbf{s}_{r})\quad,\quad\vec{h}^{\text{dst}}_{r}=\textsf{Encoder}(\mathbf{t}_{r}).$
Here we adopt the reconstruction loss $\mathcal{L}_{flow}$ to optimize the two embeddings (see Appendix A.3 for more details).
4.3.4. Region representation learning.
We aim to jointly learn region representations from spatial proximity, imagery, and flow views to extract general urban region knowledge and embed this into a unified representation space. To achieve this, we propose a view-fusion task that encodes the region embedding $\vec{h}_{r}$, the imagery embedding $\vec{h}_{r}^{img}$, and the flow embeddings $\vec{h}_{r}^{src}$ and $\vec{h}_{r}^{dst}$ into a combined vector $\hat{h}_{r}$. We then decode $\hat{h}_{r}$ using separate decoders to reconstruct the original embeddings.
To implement this idea, we employ a fusion layer to integrate the embeddings: $\hat{h}_{r}=\text{ReLU}(\textbf{W}[\vec{h}_{r},\vec{h}_{r}^{img},\vec{h}_{r}^{src},\vec{h}_{r}^{dst}]+\textbf{b})$, where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation, and W and b are trainable weights. We design a multi-view prediction task to train the fusion layer with the loss function $\mathcal{L}_{fuse}$, defined by the ability to reconstruct the representation of each individual embedding $\vec{h}_{r},\vec{h}_{r}^{img},\vec{h}_{r}^{src},\vec{h}_{r}^{dst}$ from the fused representation $\hat{h}_{r}$:
$\mathcal{L}_{fuse}=\sum_{r\in\mathcal{R}}\sum_{k=1}^{4}\|\vec{h}^{k}_{r}-\textsf{Decoder}^{k}(\hat{h}_{r})\|_{2}^{2}.$
Here, $\vec{h}^{k}_{r}$ ($k=1,2,3,4$) represents one of the four embeddings, and $\textsf{Decoder}^{k}$ is the corresponding decoder (such as a Multi-Layer Perceptron).
The final loss function can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{sp}+\mathcal{L}_{img}+\mathcal{L}_{flow}+\mu\mathcal{L}_{fuse}.$
where $\mu$ is a parameter that balances the weight of the fusion loss.
5. Prompt Task-specific Knowledge
After obtaining a general region representation, a critical issue lies in how to integrate task-specific knowledge into pre-trained representations. To achieve this, we propose two graph-based prompt tuning methods: a manually-designed prompt method based on explicit domain knowledge and a task-learnable prompt learning method to extract implicit task-specific knowledge.
5.1. Manually-designed Prompt
In the manually-designed prompt, we adjust the subgraph structure, i.e., , defining the graph pattern $\mathcal{P}(P_{V},P_{E},\theta)$ in Section 4.2, to emphasize specific relationships and entities relevant to the target task, as shown in Figure 5.
We further enhance this process by manually adjusting the proportion of different types of edges and entities in the subgraph, thereby influencing the resultant embeddings. To achieve this, we first define a set of task-specific entity weights $\mathbf{W}_{task}$ that reflect the importance of various types of entities for the downstream task. These weights are designed based on domain knowledge and the specific requirements of the task. Formally, we define $\mathbf{W}_{task}$ as:
$\mathbf{W}_{task}=\{w_{\tau_{i}}\mid\tau_{i}\in P_{V}\}.$
where $P_{V}$ is the set of all possible node types in the region-induced subgraph, and $w_{\tau_{i}}\in[0,1]$ represents the importance of entity type $\tau_{i}$. The weights $w_{\tau_{i}}$ are manually defined based on domain knowledge and are used as the deletion proportion for each entity type. We define a function $\textsf{Adjust}(\cdot)$ to delete nodes based on $\mathbf{W}_{task}$$\textsf{Adjust}(g_{r},\mathbf{W}_{task})\rightarrow g^{\prime}_{r}$ where $g_{r}=(V_{r},E_{r})$ is the original subgraph. The function Adjust obtains the adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ as follows:
(1) Node Deletion: For each entity type $\tau_{i}\in\mathcal{E}$, delete $w_{\tau_{i}}\cdot|V_{\tau_{i}}|$ nodes of type $\tau_{i}$ from $g_{r}$, where $|V_{\tau_{i}}|$ is the number of nodes of type $\tau_{i}$ in $g_{r}$, such that $V^{\prime}_{r}=V_{r}\setminus\{v\mid\phi(v)=\tau_{i}\text{ with probability }1-w_{\tau_{i}}\}$.
(2) Edge Deletion: After node deletion, delete edges connected to the removed nodes, such that $E^{\prime}_{r}=\{e\mid e=(u,v)\in E;u,v\in V^{\prime}_{r}\}$.
The adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ is then input to the pre-trained GURP to generate the task-specific embeddings.
In summary, the manually-designed prompt is the graph pattern $\mathcal{P}$ with the set of task-specific entity weights $\mathbf{W}_{task}$, which guide the $\textsf{Adjust}(\cdot)$ function in modifying the subgraph $g_{r}$ to $g^{\prime}_{r}$. By incorporating these manually-designed prompts, we can tailor the urban region representations to better capture task-relevant entities and relationships. This approach also offers the advantage of adapting to downstream tasks without the need for additional training. It simply requires adjusting the input subgraph based on the manually-defined prompts.
Figure 5. Manually-designed Prompt.
5.2. Task-learnable Prompt
Apart from leveraging explicitly designed task-specific knowledge prompts to enrich region representations, in the presence of sufficient task data, we can also learn prompts based on the task data and pre-trained representations to induce implicit knowledge within the task data, thereby enhancing region representations.
Considering that we describe the attributes of an urban region through subgraph induced with region nodes from $G$, it is intuitive to similarly describe the task adaptation to different region attributes with graphs as well. Therefore, we propose a prompt learning module to use prompt graphs to portray the adaptation of downstream task relative to the original region embedding, as shown in Figure 6.
Figure 6. Task-learnable Prompt.
We introduce the graph kernel method (Vishwanathan et al., 2010; Nikolentzos and Vazirgiannis, 2020; Feng et al., 2022) to assess the similarity of between the prompt graphs and the region subgraphs due to its excellent performance and computational efficiency, then enhancing the pre-trained region embedding. Let $\mathcal{G}$ denote the prompt graph set consisting of $m$ trainable (hidden) prompt graphs, where each prompt graph $G^{p}\in\mathcal{G}$ is defined as $G^{p}=(V^{p},E^{p},\mathbf{X}^{p})$. Here, $\mathbf{X}^{p}\in\mathbb{R}^{|V^{p}|\times d}$ represents the node attributes, with $d$ being the dimension of these attributes. For each region subgraph $g_{r}$, we utilize an attribute graph kernel function $K_{p}(\cdot,\cdot)$ to measure the similarity between $g_{r}$ and each prompt graph $G^{p}_{1},G^{p}_{2},\ldots,G^{p}_{m}$. Specifically, we employ the $P$-step random walk kernel (more details in Appendix A.4) to compute the kernel value between the region subgraph $g_{r}$ and each prompt graph $G^{p}_{i}$. These values are concatenated to form an $m$-dimensional prompt vector $\vec{h}_{p}\in\mathbb{R}^{m}$ for each region. We concatenate the prompt vector $\vec{h}_{p}\in\mathbb{R}^{m}$ before the pre-trained region representation $\vec{h}_{r}$ to derive region embedding $\vec{h}^{\prime}_{r}=\vec{h}_{p}||\vec{h}_{r}$. The new embeddings are then fed into a linear layer $\textsf{Linear}(\cdot)$. We update the prompt graphs and linear layer parameters, by optimizing on the downstream labeled task, to obtain the final prompt representation with the simple MSE loss function:
$\mathcal{L}_{task}=\frac{1}{N}\sum_{i=1}^{N}(y^{pred}_{i}-y^{label}_{i})^{2}.$
6. Experiments
We conduct extensive experiments to investigate the following research questions (RQs):
  • RQ1: Can our proposed pre-train and prompt framework (GURPP) outperform previous approaches?
  • RQ2: Does GURP have the ability to capture and apply heterogeneous and generic patterns?
  • RQ3: How does the manually-designed prompt work?
  • RQ4: How does the task-learnable prompt perform and how does it look like?
  • RQ5: How does each component contribute to GURPP?
6.1. Experimental Setup
6.1.1. Datasets.
We collect the datasets of region division, POI, urban imagery, taxi trips, road network, check-in, crime and crash for two representative cities in the United States: New York City (NYC) and Chicago (CHI). The detailed statistics and sources of the dataset is available in Appendix A.5.
6.1.2. Baselines.
We compare our model with the following representative baseline models: knowledge graph embedding methods and its variants (TransR (Lin et al., 2015), TransR-N, TransR-G), graph embedding methods (node2vec (Grover and Leskovec, 2016), GAE (Kipf and Welling, 2016)), and urban region embedding methods (MVURE (Zhang et al., 2021), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023a), HRE, ReCP (Li et al., 2024)). These baselines cover the range from traditional graph embedding to complex models that take into account urban-specific attributes, providing us with a comprehensive performance comparison benchmark. Specific details and descriptions can be found in the Appendix A.6.
6.1.3. Metrics and Implementation.
To evaluate the prediction performance, we use three widely recognized evaluation metrics: the coefficient of determination ($R^{2}$), root mean squared error (RMSE), and mean absolute error (MAE). Better performance is indicated by a higher $R^{2}$ and lower RMSE and MAE values. Our model includes three version: 1) GURP represents the pre-train module, 2) $\text{GURPP}_{M}$ is GURP with manually-designed prompt, and 3) $\text{GURPP}_{T}$ is GURP with task-learnable prompt. For detailed implementation, please refer to Appendix A.7 and our provided code repository.
Table 1. Main performance comparison of different models on downstream tasks across NYC and CHI datasets. The best results for each metric are highlighted in bold, the second-best are underlined, and the third-best are marked with an asterisk. The last and the third-to-last rows show the percentage improvement of the GURPP and GURP models over the third-best model.
Model New York City (NYC) Chicago (CHI)
Crime Prediction Check-in Prediction Crime Prediction Crash Prediction
MAE RSME $R^{2}$ MAE RSME $R^{2}$ MAE RSME $R^{2}$ MAE RSME $R^{2}$
TransR-N 111.319 148.863 -0.014 534.699 805.282 0.031 209.072 307.454 -0.037 106.914 176.860 0.001
TransR-G 109.221 146.905 0.013 530.427 802.426 0.038 208.404 306.155 -0.028 104.680 174.419 0.028
node2vec 138.458 175.797 -0.414 573.118 837.441 -0.048 246.066 349.965 -0.343 121.900 174.907 0.023
GAE 96.846 126.364 0.270 349.942 542.920 0.560 200.606* 300.237 0.011 100.536 173.323 0.040
MVURE 72.471* 99.844 0.544 310.973 498.076 0.629 206.080 289.859* 0.079* 96.953 148.865* 0.292*
MGFN 73.502 97.440* 0.566* 300.150 456.626* 0.688* 214.021 310.832 -0.059 108.3740 179.854 -0.033
HRE 76.926 102.07 0.523 329.547 513.801 0.606 206.612 298.389 0.024 99.952 158.625 0.196
HREP 74.854 100.035 0.542 320.335 505.516 0.618 206.591 298.238 0.025 99.766 158.658 0.196
ReCP 73.853 100.558 0.537 292.871* 524.890 0.588 207.485 298.965 0.020 93.298* 157.515 0.207
GURP 72.094 93.216 0.603 271.689 412.918 0.745 161.266 225.244 0.444 80.362 108.781 0.622
Enhance 0.52% 4.33% 6.54% 7.23% 9.57% 8.28% 19.61% 22.29% 462.03% 13.87% 26.93% 113.01%
GURPP 71.481 91.612 0.616 261.450 388.225 0.775 146.385 203.050 0.548 65.297 92.114 0.729
Enhance 1.36% 5.98% 8.83% 10.73% 14.98% 12.65% 27.03% 29.95% 593.67% 30.01% 38.12% 149.66%
6.2. RQ1 & RQ2: Overall Performance
We evaluate the overall performance of baselines and our model on three downstream tasks. The results of the experiments are shown in Table 1, where the GURPP is all implemented by $\text{GURPP}_{T}$. From the results, we find the following:
(1) GURPP demonstrates superior performance to baselines in both pre-training and prompting phases. In the pre-training phase, the performance of GURP improves by 6.54%, 8.28%, 462.03%, and 113.01% over the best baseline of $R^{2}$ on two urban datasets. Meanwhile, on all downstream tasks, the average performance on MAE and RMSE metrics improves by 17.28% and 22.26%, respectively. Subsequently, the performance of GURPP is further improved by an average of 9.50%, 9.28%, and 11.703% of MAE, RMSE and $R^{2}$, respectively, when we introduce the task-learnable prompt. The results validate the significant advantages of our proposed pre-train and prompt framework.
(2) To validate the ability of our pre-training method to capture generic urban knowledge, we compare the performance of state-of-the-art urban region embedding methods on datasets from two distinct cities, NYC and CHI. The results show that while these baselines perform well on the NYC dataset, their performance drops significantly on the CHI dataset. For instance, the performance on CHI decreases by an average of 71.8% in $R^{2}$ compared to NYC. In contrast, our GURPP model demonstrates remarkable consistency, with only an 8.49% decrease in $R^{2}$. This may be due to the fact that the baselines tend to be effective only at extracting information specific to certain data types, but they do not consider spatial structure and functional layouts among urban entities. For example, existing baselines rely on the flow view to extract information. However, the content of flow view in the CHI dataset is sparse, leading to a significant degradation in the performance of downstream task. While GURPP shows excellent performance on both urban datasets, which demonstrates the robustness and universality of GURPP across diverse urban environments.
(3) To further validate the effectiveness of our subgraph-centric and multi-view fusion module, we conduct experiments comparing the performance of models with and without subgraphs and multi-view fusion. Specifically, we compare the TransR-G model (adopts region-induced subgraph and regards graph-level embedding as the region embedding) with the TransR-N model (directly uses region nodes embedding as region embedding). The results indicate that TransR-G outperforms TransR-N by an average of 1.27% in MAE, underscoring the effectiveness of the subgraph-centered approach. This improvement is attributed to the subgraph-centric method’s ability to better preserve the structural and contextual information of regions, providing more accurate and meaningful urban region embeddings. The average 32.55% improvement in MAE for GURP over TransR-G is mainly due to the multi-view fusion strategy, which improves the model’s comprehensive understanding of urban features.
6.3. RQ3: Manually-designed Prompt Study
In the real world, the incidence of crime is often closely linked to specific POI categories, roads and junctions, for example, ATMs, bars, parks, etc. are more likely to be the location of a crime. The density of roads and the complexity of junctions are also important factors affecting crime rates. As for the check-in behavior, the brand of the POI is given more attention, and certain users may be more inclined to check in at locations of well-known brands such as Starbucks, McDonald’s, or other specific brands. Therefore, we designed two patterns as shown in Figure 7 to validate the performance of manually-designed prompt.
The results are shown in Table 2, from which it can be found that crime prediction does pay more attention to factors such as POI categories, roads and junctions, which are closely related to criminal behavior. Meanwhile, removing POI brand information shows a significant decrease in the performance of check-in prediction. These results also demonstrate that the prediction performance of the model can be improved by manually-designed prompt.
Figure 7. The structure of the manually-designed prompt based on Pattern 1 (P1) and Pattern 2 (P2). P1 ignores POI brand information. P2 randomly samples POI, road, and junction nodes around the region with a weight of 0.9.
Table 2. Performance of $\text{GURPP}_{M}$.
Model Crime Prediction (NYC) Check-in Prediction (NYC)
MAE RSME $R^{2}$ MAE RSME $R^{2}$
GURP 72.094 93.216 0.603 271.689 412.918 0.745
$\text{GURPP}_{M}$-P1 68.538 88.698 0.640 273.201 429.564 0.724
$\text{GURPP}_{M}$-P2 69.421 91.834 0.614 262.437 397.466 0.764
6.4. RQ4: Task-learnable Prompt Study
6.4.1. Effectiveness of Task-learnable Prompt.
We utilize GURP pre-trained embedding with the simple prefix prompt (proposed by HREP(Zhou et al., 2023a)) and task-learnable prompt respectively for comparative analysis. Both of these prompt methods are data-driven and seek to improve the region representation by learning data for specific downstream tasks. We conduct experiments on crime prediction task in NYC and crash prediction task in CHI, respectively, and the results are shown in Table 3.
Our task-learnable prompt outperforms pre-trained embeddings across both tasks, with improvements of 8.83% of $R^{2}$ in crime prediction and 149.66% of $R^{2}$ in crash prediction, indicating that the prompt is highly effective in providing tailored guidelines for different downstream tasks. However, in crime prediction task, the prefix prompt have an negative effect, resulting in a performance decrease of 1.16% of $R^{2}$. This demonstrates that our task-learnable prompt is more stable and effective compared to prefix prompt.
Table 3. Effectiveness of Task-learnable Prompt.
Framework Crime Prediction (NYC) Crash Prediction (CHI)
MAE RSME $R^{2}$ MAE RSME $R^{2}$
GURP 72.094 93.216 0.603 80.362 108.781 0.622
GURP+Pre 73.758 93.971 0.596 66.330 93.772 0.719
$\text{GURPP}_{T}$ 71.481 91.612 0.616 65.297 92.114 0.729
6.4.2. Visualization of Learned Prompt Graph.
Compared to prefix prompt, our task-learnable prompt is able to generate a series of prompt graphs during training, which enhances the clarity of our prompt method. We visualize prompt graphs that reveal key structural features in the urban region graph of interest in the task. In Figure 8, we show some prompt graphs for crime prediction tasks in CHI. (Additional prompt graphs can be found in Appendix A.8.) These results show that the prompt graphs generally exhibits circular structural features, implying that the model have identified and emphasized interactions between urban entities.
Figure 8. Visualization of learned prompt graphs.
6.5. RQ5: Ablation Studies
Finally, we conduct the ablation studies in NYC dataset to investigate the effectiveness of the individual components of GURPP. Variants of GURPP are designed as follows: GURPP/P. We use only the pre-training process to learn urban region general representation.GURPP/F,  GURPP/S,  GURPP/I. We remove the flow view, spatial proximity view, and imagery view from GURPP, respectively.  GURPP/K. We replace the knowledge graph embedding initialization with random initialization.
Figure 9 presents the ablation results of our model and its variants. Spatial proximity view and knowledge graph embedding method significantly enhance the model’s performance, lifting it by 151.64% and 156.67% of $R^{2}$, respectively. These components effectively capture the heterogeneous and generic patterns of interactions among entities by elucidating the spatial relationships and structured knowledge of the urban region. The flow view and the imagery view can also greatly enhance the performance of the model, increasing it by 29.04% and 21.11%, respectively. This is because they capture dynamic flow information and visual features of urban regions, respectively, providing the model with a multi-dimensional context that enhances its ability to fully understand and predict urban dynamics. Prompt learning, with an improvement of 3.12%, effectively elevates model performance by introducing guiding prompts that allow the model to delve deeper into understanding and utilizing task-relevant knowledge.
Figure 9. Results of ablation studies.
7. Related Work
Extensive studies have been conducted to learn urban region representation from different perspectives. We combine the existing work categories in Table 4.
Table 4. Cats.of Related Works on Urban Region Embedding
Methods Task-specific learning or fine-tuning Prompting
Non-graph (Yang et al., 2022; Zhou et al., 2023b; Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Jenkins et al., 2019) (Yuan et al., 2024; Yan et al., 2024; Zhou et al., 2023a)
Graph (Xu et al., 2023; Zhang et al., 2024; Bai et al., 2019; Han et al., 2020; Zhang et al., 2021; Wu et al., 2022; Liu et al., 2023a; Ning et al., 2024) GURPP(Ours)
Task-specific learning or fine-tuning. Urban region representation learning aims to extract complex patterns and relationships within urban regions and embed them into a representation space. Existing research can be broadly classified into two categories: task-specific supervised region representation learning (Bai et al., 2019; Han et al., 2020; Yang et al., 2022; Zhou et al., 2023b) and general unsupervised region representation learning (Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021). These works utilize feature aggregation models to generate representation from predefined region statistics with limitations in capturing detailed spatial layout and semantics. Moreover, using predefined statistical features makes it challenging to further adapt the representation to the specificities of various downstream tasks. We address the limitations by leveraging a urban region graph and subgraph-centric region embeddings, which learn the generic knowledge of urban spatial layout and generate flexible region representation tailored to different tasks.
Pre-training and Prompting. Recent years have witnessed the significant success of pre-training and prompting paradigm in fields of NLP (Liu et al., 2023b; Brown et al., 2020; Roberts et al., 2020; Gao et al., 2021), CV (Yoo et al., 2023; Radford et al., 2021; Zhou et al., 2022b, a) and graph learning (Sun et al., 2023; Yu et al., 2024b; Ma et al., 2024; Yu et al., 2024a). Many works in urban computing have also made attempts in this regard (Xu et al., 2023; Zhang et al., 2024; Balsebre et al., 2023; Yuan et al., 2024; Yan et al., 2024). For instance, Yuan et al. (Yuan et al., 2024) design an urban spatio-temporal prediction pre-train and prompt model. Yan et al. (Yan et al., 2024) introduce large language models (LLMs) to describe urban imagery features and unify different prediction tasks through textual prompts. However, the urban tasks explored in these work differ from ours, with a distinct focus. Our work can also benefit from the integration of LLMs to enhance fine-grained characterization capabilities. As a similar work, HREP (Zhou et al., 2023a) introduces prefix prompt to region embedding, but the pre-trained model fails to adequately capture the fine-grained features of regions. We propose both manually-designed and task-learnable prompts built on graph-based pre-train model to address this issue.
8. Conclusion and Future Works
In this paper, we introduce the urban region graph and propose a graph-based pre-training and prompting framework for urban region representation learning, balancing urban generality and task specificity. We develop a subgraph-centric pre-training model to capture general urban knowledge and two prompt methods to incorporate task-specific insights. Extensive experiments demonstrate the effectiveness of our approach. Future work may explore: 1) Integrating textual data and LLMs to uncover more universal urban knowledge; 2) Developing arbitrary region representations for flexible urban modeling; 3) Assessing the transferability of urban representations across cities to enhance general applicability.
Appendix A Appendix
A.1. Appendix of urban region graph construction
Here we provide more details for the urban region graph construction. Specifically, we incorporate brand information for POI entities, such as Starbucks, to enhance the representation of POI functionalities. Altogether, we define eight node types shared across regions and cities: region, POI, POI category, brand, road, road category, junction, and junction category, comprising the node type set $T_{V}$. We also define six edge types, forming the edge type set $T_{E}$. These include the NearBy relationship, which connects region entities based on their geographical adjacency; the Contains relationship, which is based on spatial containment; the BrandOf relationship, which links brands to POI entities; and the CateOf, JCateOf, and RCateOf relationships, which are based on the category information of the entities.
A.2. Appendix of region imagery encoding
We supplement more details for the urban imagery encoding. Specifically, we employ ResNet (He et al., 2016) to extract visual features for each region image, such that $\vec{a}_{k}=\text{ResNet}(a_{k})$ and derive the comprehensive visual representation for the region by averaging the features of multiple images within the region. Then, a linear projection layer is utilized to project the visual feature into the representation space of region attribute, such that
$\vec{h}_{r}^{img}=\text{Linear}(\frac{1}{m}\sum_{k=1}^{m}\vec{a}_{k}).$
The contrastive loss $\mathcal{L}_{img}$ is then used to align the visual representation with the subgraph-centric region representation, ensuring a cohesive integration of visual and attribute information.
$\mathcal{L}_{img}$ $=\sum_{r\in\mathcal{R}}\Big{[}-\log\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)$
$+\log\left(\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)+\sum_{j=1}^{B-1}\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{j}}{\tau}\right)\right)\Big{]}$
Here, $B$ represents batch size and $\tau$ is the temperature parameter.
A.3. Appendix of human mobility encoding
Here we give a detailed description for human mobility encoding. To learn the embeddings, we reconstruct the departure and arrival flow distribution. Let $m_{ij}$ represent the number of trips from region $r_{i}$ to region $r_{j}$. The source and destination urban flow distributions are computed as $P^{{src}}(r_{j}|r_{i})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{ik}}$ and $P^{{dst}}(r_{i}|r_{j})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{kj}}$, respectively. We then reconstruct the source and destination distributions $\hat{P}^{{src}}(r_{j}|r_{i})$ and $\hat{P}^{{dst}}(r_{i}|r_{j})$ using the encoded inflow and outflow features:
$\hat{P}^{{src}}(r_{j}|r_{i})=\frac{\exp((\vec{h}_{i}^{{src}})^{\top}\vec{h}_{j}^{{dst}})}{\sum_{k}\exp((\vec{h}_{i}^{{src}})^{\top}\vec{h}_{k}^{{dst}})}$
$\hat{P}^{{dst}}(r_{i}|r_{j})=\frac{\exp((\vec{h}_{j}^{{dst}})^{\top}\vec{h}_{i}^{{src}})}{\sum_{k}\exp((\vec{h}_{j}^{{dst}})^{\top}\vec{h}_{k}^{{src}})}$
Finally, we learn the inflow and outflow embeddings by optimizing the following loss function:
$\mathcal{L}_{flow}=\sum_{(r_{i},r_{j})\in\mathcal{R}\times\mathcal{R}}$ $-P^{src}(r_{j}|r_{i})\log\hat{P}^{src}(r_{j}|r_{i})$
$-P^{dst}(r_{i}|r_{j})\log\hat{P}^{dst}(r_{i}|r_{j}),$
which helps in accurately learning the inflow and outflow embeddings by comparing the true distributions with the reconstructed distributions.
A.4. Appendix of attribute graph kernel function
We introduce how attribute graph kernel function $K_{p}(\cdot,\cdot)$ works. We first define the graph kernel function $K(\cdot,\cdot)$. The graph kernel function is a $P$-step random walk kernel (Nikolentzos and Vazirgiannis, 2020) to measure the similarity between graphs, which counts the number of common walks in both graphs within $P$ steps with its calculation as follows. Given $G_{1}=(V_{1},E_{1}),~{}G_{2}=(V_{2},E_{2})$, where $V$ and $E$ are the set of nodes and edges, respectively, we initially define the direct product graph $G_{\times}$ following the framework used in existing research (Nikolentzos and Vazirgiannis, 2020; Feng et al., 2022):
(1) $G_{\times}=G_{1}\times G_{2}=(V_{\times},E_{\times})$
where $V_{\times}=\{(v_{1},v_{2}):v_{1}\in V_{1},v_{2}\in V_{2}\}$, $E_{\times}=\{\{(v_{1},v_{2}),(u_{1},u_{2})\}:\{v_{1},u_{1}\}\in E_{1}\land\{v_{2},u_{2}\}\in E_{2}\}$. Following the theorem (Vishwanathan et al., 2010) that a single random walk on the direct product graph $G_{\times}$ is equivalent to a random walk on the two graphs $G_{1},~{}G_{2}$ at the same time, the $P$-step random walk kernel is computed in the following way:
(2) $K(G_{1},G_{2})=\sum_{p=0}^{P}K_{p}(G_{1},G_{2})=\sum_{p=0}^{P}\lambda_{p}\sum_{i,j=1}^{|V_{\times}|}[A_{\times}^{p}]_{ij}$
where $A_{\times}$ is the adjacency matrix of the direct product graph, $\lambda_{p}$ is the sequence of weights and the $(i,j)$ item of $A_{\times}^{p}$ represents the $P$-length common walks between the $i$-th and $j$-th nodes.
Extending the above definition to the scenario of graphs with node attributes, we use $\mathbf{X}\in\mathbb{R}^{n\times d}$ to denote the attribute matrix of a graph with $n$ nodes and attribute dimension $d$. Specifically, the node attribute matrices of $G_{1},G_{2}$ are denoted by $\mathbf{X}_{1}\in\mathbb{R}^{n_{1}\times d},\mathbf{X}_{2}\in\mathbb{R}^{n_{2}\times d}$. Thus, for the graph $G_{\times}=G_{1}\times G_{2}$, the node attribute matrix is $\mathbf{S}=\mathbf{X}_{1}\mathbf{X}_{2}^{\top},\mathbf{S}\in\mathbb{R}^{n_{1}\times n_{2}}$, where the item $(i,j)$ encodes the similarity between the $i$-th node in $G_{1}$ and the $j$-th node in $G_{2}$. We flattened $\mathbf{S}$ to $\mathbf{s}\in\mathbb{R}^{n_{1}n_{2}}$ to obtain the formula for the random walk kernel:
(3) $K_{p}(G_{1},G_{2})=\sum_{i,j=1}^{|V_{\times}|}\mathbf{s}_{i}\mathbf{s}_{j}[A_{\times}^{p}]_{ij}=\mathbf{s}^{\top}A^{p}_{\times}\mathbf{s}.$
A.5. Appendix of dataset statistics and sources
We collect the datasets of region division, POI, urban imagery, taxi trips, road network, check-in, crime and crash for two representative cities in the United States: New York City (NYC) and Chicago (CHI). The region division is based on census tracts and the urban imagery are 256 x 256 pixels with a resolution of 4.7 meters. The detailed statistics and sources of the dataset is shown in Table 5.
Table 5. Dataset Statistics and Sources.
Data NYC CHI Source
Region 180 869 Census Bureau (Census Bureau, 2024)
POI 143,445 112,989 OSM (OpenStreetMap, 2024), Safegraph (Safegraph, 2024)
Imagery 7,160 - ArcGIS (Arcgis, 2024), UUKG (Liu et al., 2023c)
Taxi trip 12,684,353 3,889,032 NYCOD (NYC OpenData, 2024), CHIDP (Chicago Data Portal, 2024)
Road 110,919 139,850 OSM (OpenStreetMap, 2024)
Junction 62,437 66,065 OSM (OpenStreetMap, 2024)
Crime 35,335 274,733 NYCOD (NYC OpenData, 2024), CHIDP (Chicago Data Portal, 2024)
Check-in 106,902 - Foursquare (Yang et al., 2014; Foursquare, 2024)
Crash - 150,535 CHIDP (Chicago Data Portal, 2024)
A.6. Appendix of baselines description
We compare our model with representative baseline models, including graph embedding methods, knowledge graph embedding methods and state-of-the-art region embedding methods.
I. Knowledge graph embedding methods:
  • TransR (Lin et al., 2015) captures the complex interactions between entities and relationships in the knowledge graph by projecting them into different spaces and learning their representation in these spaces.
  • TransR-N method uses the embedding of the region node learned by TransR as the region representation.
  • TransR-G method is to average the embedding of the region subgraph learned by TransR as the region representation.
II. Graph embedding methods:
  • node2vec (Grover and Leskovec, 2016) generates a sequence of nodes by performing random walks on the graph to capture the neighborhood information of nodes, and then learns the representation of nodes using the Skip-Gram model.
  • GAE (Kipf and Welling, 2016) consists of an encoder and a decoder. The encoder is used to learn the representation of nodes, and the decoder tries to reconstruct the adjacency matrix from these node representation.
III. State-of-the-art urban region embedding methods:
  • MVURE (Zhang et al., 2021) constructs intra- and inter-region data multi-view based on human mobility and region attributes, and then performs multi-view fusion to learn region representation.
  • MGFN (Wu et al., 2022) builds human mobility patterns and then learns intra- and inter-pattern relationships to obtain region representation through a multi-level cross-attention mechanism.
  • HREP (Zhou et al., 2023a) designs a multi-relational heterogeneous graph of regions and used a relation-aware GCN to learn region representation, as well as a prefix prompt mechanism for downstream tasks.
  • HRE is a part of HREP without prompt learning.
  • ReCP (Li et al., 2024) is a multi-view comparative prediction model that learns region representation by capturing the unique information of a single view as well as the consistency between two views.
A.7. Appendix of Implementation
In our experiments, all the region embedding dimensions is set to 144. The number of HGT layer and head is 2 and 4, and the triplet margin parameter $\delta=2$, fusion loss weight $\mu=0.01$. We set the learning rate to 0.001 and the weight decay to 1e-6. In the task-learnable prompt module, we set the number of random walk kernel to 2 and the size of prompt graph to 6 and 8 of each layer.
We implemented our model with Pytorch 1.8, and conducted experiments on a machine equipped with a NVIDIA RTX 3090 GPU, a 64 GB RAM and a 3.50 GHz CPU.
A.8. Appendix of prompt graphs
Figure 10 displays additional learned prompt graphs, with the majority of them being circular structural features.
Figure 10. Visualization of additional learned prompt graphs.

### v4

@@
 However, despite the proliferation of methods and their success, acquiring general urban region knowledge and adapting to different tasks remains challenging.
-Previous work often neglects the spatial structures and functional layouts between entities, limiting their ability to capture transferable knowledge across regions.
-Further, these methods struggle to adapt effectively to specific downstream tasks, as they do not adequately address the unique features and relationships required for different downstream tasks.
+Existing work pays limited attention to the fine-grained functional layout semantics in urban regions, limiting their ability to capture transferable knowledge across regions.
+Further, inadequate handling of the unique features and relationships required for different downstream tasks may also hinder effective task adaptation.
 In this paper, we propose a Graph-based Urban Region Pre-training and Prompting framework (GURPP) for region representation learning.
-Specifically, we first construct an urban region graph that integrates detailed spatial entity data for more effective urban region representation.
-Then, we develop a subgraph-centric urban region pre-training model to capture the heterogeneous and transferable patterns of interactions among entities.
-To further enhance the adaptability of these embeddings to different tasks, we design two graph-based prompting methods to incorporate explicit/hidden task knowledge.
-Extensive experiments on various urban region prediction tasks and different cities demonstrate the superior performance of our GURPP framework.
-The implementation is available at this repository: https://anonymous.4open.science/r/GURPP.
-Urban Computing, Urban Region Representation Learning
-copyright: acmlicensed
+Specifically, we first construct an urban region graph and develop a subgraph-centric urban region pre-training model to capture the heterogeneous and transferable patterns of entity interactions.
+This model pre-trains knowledge-rich region embeddings using contrastive learning and multi-view learning methods.
+To further refine these representations, we design two graph-based prompting methods: a manually-defined prompt to incorporate explicit task knowledge and a task-learnable prompt to discover hidden knowledge, which enhances the adaptability of these embeddings to different tasks.
+Extensive experiments on various urban region prediction tasks and different cities demonstrate the superior performance of our framework.
+Urban Foundation Model, Region Representation, Graph Prompt
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadadoi: XX.XXXX/XXXXXXX.XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XXXX/XXccs: Information systems Data miningccs: Computing methodologies Knowledge representation and reasoning
 1. Introduction
-An urban region is a geographical area comprising diverse spatial entities (e.g.,  shopping centers, restaurants, and roads) and their interactions (e.g.,  nearby and located in).
+An urban region is a geographical area comprising diverse spatial entities (e.g.,  shopping centers and roads) and their interactions (e.g.,  nearby and located in).
 Recently, urban region representation learning, which aims to extract vector embeddings from these entities, has become a significant research focus (Zhang et al., 2024).
-Effective urban region representations can be applied to various tasks, such as region popularity prediction (Li et al., 2024; Fu et al., 2019; Zhou et al., 2023a; Wu et al., 2022), house price prediction (Wang and Li, 2017), crime prediction (Wang and Li, 2017; Zhou et al., 2023a; Wu et al., 2022; Zhang et al., 2021), and population density inference (Li et al., 2023), enabling urban planners, policymakers, and businesses to enhance decision-making processes.
+Effective urban region representations can be applied to various tasks, such as region popularity prediction (Li et al., 2024a; Fu et al., 2019; Zhou et al., 2023a; Wu et al., 2022), house price prediction (Wang and Li, 2017), and crime prediction (Wang and Li, 2017; Zhou et al., 2023a; Wu et al., 2022; Zhang et al., 2021), enabling urban planners and businesses to enhance decision-making processes.
+Figure 1. Example of urban generality and task specificity.
 However, the complexity of urban environments and the diversity of applications pose significant challenges.
-Figure 1. Example of urban generality and task specificity.
-Balancing urban generality and task specificity is crucial in urban region representation.
-(i) Urban regions often appear similar in their spatial structures and functional layouts, because the interactions among entities within these regions follow similar patterns.
+On the one hand, urban regions often appear similar in their spatial structures and functional layouts, because the interactions among entities within these regions follow similar patterns.
 These shared patterns serve as generic knowledge across urban areas, facilitating the development of generalizable region representation models.
-As illustrated in Figure 1, Region A and Region B showcase typical urban regions.
+As illustrated in Figure 1, Regions A and B showcase typical urban regions.
 In these examples, spatial structures are demonstrated by the proximity of residences to parks and shops, which enhances community convenience and accessibility.
@@
 Therefore, an effective urban region representation model needs to capture the heterogeneous and generic patterns of interactions among entities within regions.
-(ii) Despite the broad similarities, each urban region and downstream task possesses unique details.
+On the other hand, despite the broad similarities, each urban region and downstream task possesses unique details.
 Specific tasks may require attention to different features within urban data.
-For example, in traffic crash prediction, the density of roads and the number of residents living nearby, might be critical features, as they directly influence the probability of accidents occurring.
-On the other hand, crime prediction models might focus on the density and type of POIs within the region, such as bars or late-night convenience stores.
-As consequence, balance between generality and specificity raises a crucial question: Can we design an urban region representation framework that effectively integrates generalizable knowledge while accommodating the specific needs of distinct tasks?
-Recently, the pre-training and prompting paradigm that learns general representations that can be fine-tuned or prompted to adapt to specific tasks has shown remarkable success in the field of natural language processing (Liu et al., 2023b; Brown et al., 2020; Roberts et al., 2020; Gao et al., 2021) and computer vision (Radford et al., 2021; Zhou et al., 2022b, a).
-Inspired by these advancements, we hope to design a pre-training and prompting framework for urban region representation learning.
-However, building such a framework for urban representation is never easy.
-The challenges are as follows.
-(1) Capturing the heterogeneous and generic patterns of interactions among spatial entities within regions is hard.
-Most previous methods follow a process where they first predefine aggregations and statistics about urban entities and relationships, and then learn from these predefined features (Han et al., 2020; Yang et al., 2022; Zhang et al., 2023; Li et al., 2023).
-While this approach has its merits, it often fails to capture the spatial structures and functional layouts between entities within regions which is generalizable.
-Other previous methods consider urban entities in isolation, but they focus on only a few types, such as POIs and roads and a few relations between them.
-This narrow focus limits their ability to capture the heterogeneous and generic patterns of interactions among entities, resulting in embeddings that are less robust and less effective (Fu et al., 2019).
-(2) Tuning pre-trained region embeddings to capture the tasks’ focus on different features or relationships within the original urban data poses a challenge.
-The majority of previous urban region representation methods use simple regression models which merely projects the pre-train embeddings without capturing the specific entities and relationships relevant to individual tasks (Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021).
-In contrast, prompting offers a more flexible approach by allowing models to focus on specific aspects of the data as needed (Liu et al., 2023b; Yoo et al., 2023; Sun et al., 2023), effectively highlighting small but crucial differences between urban downstream tasks.
-Despite this potential, existing prompting methods (Yan et al., 2024; Zhou et al., 2023a) have limitations.
-These methods use simple prompts that do not consider detailed and heterogeneous interactions among various urban entities.
-Additionally, many existing prompting techniques simply concatenate prompts onto the original embeddings, lacking the flexibility to selectively filter out entities and relationships that are specifically useful for downstream tasks.
-This limitation makes it challenging to tailor the generalizable embedding precisely to the unique characteristics of different urban downstream tasks.
-Figure 2. Main idea of urban pre-training and prompting
-In light of the challenges discussed, we propose a novel framework named GURPP (Graph-based Urban Region Pre-training and Prompting Framework) as shown in Figure 2. Specifically:
-To address the first challenge, we develop the Subgraph-centric Urban Region Pre-training (GURP) model.
-To capture the spatial structures and functional layouts between entities, we construct an urban region graph that integrates urban entities and their relationships.
-Unlike previous methods where each region is treated as a single point, our approach views each region as a subgraph.
-So the GURP model then extract region subgraphs from urban region graph with graph patterns and learns region embeddings from these subgraphs through a heterogeneous graph learning model.
-To further enhance the learning process, we design a learning pipeline that incorporates heterogeneous graph learning model and multi-view learning.
-It enables the model to learn both structure information within subgraphs and the relationships between subgraphs.
-To address the second challenge, we develop a knowledge-enriched prompting learning model.
-This model includes two approaches: a manually-designed prompt method and a task-learnable prompt learning method.
-The manually-designed prompt adjusts subgraph structures to highlight task-specific relationships and entities using predefined weights.
-The task-learnable prompt learns from task data to generate prompt graphs that capture implicit task-specific knowledge.
-By incorporating both predefined and learned task-specific prompts, the model tailors the general embeddings to better suit the unique characteristics and requirements of various urban computing tasks.
-Contributions:
-  • (Urban region graph). We introduce urban region graph, which integrates detailed spatial and entity data for more effective urban region representation.
-  • (Pre-training model).
-We propose a Subgraph-centric Urban Region Pre-training (GURP) model for learning embeddings from urban region graphs, capturing generalizable patterns of interactions among entities within regions.
-  • (Prompting model).
-We design a novel knowledge-enriched prompting learning model with two graph-based prompting methods: a manually-defined prompt to incorporate explicit task knowledge and a task-learnable prompt to discover hidden task-relevant patterns.
-  • (Extensive experiments).
-We demonstrate the effectiveness of our GURPP framework through extensive experiments on three urban region prediction tasks across two cities, achieving significant performance improvements with an average increase of $17.28\%$ on MAE and $22.26\%$ on RMSE.
-2. Preliminaries
+For example, in traffic crash prediction, the density of roads and the number of residents living nearby might be critical features, as they directly influence the probability of accidents occurring.
+Meanwhile, crime prediction models might focus on the density and type of POIs (Point-of-Interests) within the region, such as bars or late-night convenience stores.
+As a consequence, the balance between generality and specificity raises a crucial question: Can we design an urban region representation framework that effectively integrates generalizable knowledge while accommodating the specific needs of distinct tasks?
+To solve this problem, this paper presents a simple yet effective framework with pre-training and prompting on geographical data, since this new paradigm has shown remarkable success in balancing embedding generality and task specificity in the areas like natural language processing (Liu et al., 2023c; Roberts et al., 2020), computer vision (Zhou et al., 2022b, a), and graphs (Sun et al., 2023, 2024b).
+The intrinsic reason behind this superiority, as many researchers reported (Wang et al., 2025), lies in the powerful capability of data manipulation from the prompt side, and knowledge learning from the pre-training side, which can offer flexible adaptation for various downstream tasks and useful knowledge for generalizable representation.
+However, handling such complex non-linear geographical data requires prior experience.
+Challenges are twofold:
+  • $\circ$
+    Difficulties in capturing the heterogeneous and generic knowledge from complex spatial entity interactions.
+Existing methods mostly predefine aggregations and statistics of some urban entities and relationships, and then use these these predefined features for downstream tasks (Yang et al., 2022; Zhang et al., 2023; Li et al., 2023), which often fail to capture the spatial structures and functional layouts between various entities within regions.
+  • $\circ$
+    Difficulties in capturing task-specific patterns.
+Existing methods usually use simple prompts like directly concatenating them to the pre-trained embeddings (Yan et al., 2024; Zhou et al., 2023a), which might work for certain domains but might fall short for urban region tasks with heterogeneous and complex geographical entity interactions.
+Figure 2. Main idea of urban pre-training and prompting.
+In response to these challenges, we propose GURPP, a Graph-based Urban Region Pre-training and Prompting Framework as shown in Figure 2.
+Specifically, GURPP leverages the expressive power of urban graphs to capture both generic and task-specific urban features.
+By constructing the urban region graph, GURPP effectively extracts the common spatial structures and functional layouts inherent to urban areas.
+Simultaneously, GURPP addresses the specificity required for different downstream tasks through a two-stage process.
+First, a subgraph-centric pre-training model, enhanced by multi-view self-supervised learning, captures diverse urban patterns—ranging from mobility trends and region imagery to intricate spatial and functional relationships.
+Then, two complementary graph-based prompting strategies further refine these representations: a manually-designed prompt that incorporates explicit task-specific semantics by emphasizing critical relationships and entities, and a task-learnable prompt that adapts to the hidden complexities of urban data.
+This design aims to bridge the gap between general urban knowledge and task-specific requirements for urban region representation.
+Contributions are as follows.
+  • $\circ$
+    Graph-based Urban Region Framework: We introduce GURPP, a novel graph-based pre-training & prompting framework for urban region representation, which unifies the modeling of complex spatial entity semantics.
+This design enables the extraction of heterogeneous and generic urban features, overcoming the limitations of methods that rely on predefined aggregations.
+  • $\circ$
+    Subgraph-centric Pre-training and Dual Prompting Models: We develop a novel subgraph-centric pre-training model, integrated with multi-view self-supervised learning, to capture urban generic knowledge from various modalities.
+Additionally, we propose two graph-based prompting methods—a manually-designed prompt and a task-learnable prompt—to inject explicit and implicit task-specific signals into the representations, thereby enhancing adaptability to diverse urban prediction tasks.
+  • $\circ$
+    Extensive Empirical Validation: We demonstrate the effectiveness of GURPP through extensive experiments on six urban region prediction tasks across two cities.
+Our results show significant performance improvements, with an average increase of 13.57% on MAE and 16.24% on RMSE, highlighting the model’s strong capabilities in real-world urban prediction tasks.
+Figure 3.
+Overview of GURPP framework.
+(1) Construct the urban region graph that integrates specific urban entities and semantic relationships.
+(2) Extract a subgraph of each region according to the graph pattern.
+(3) Learn the subgraph-enhanced region representations with multi-view fusion.
+(4) Adapt the region embeddings with a manually-designed prompt method and a task-learnable prompt learning method.
+2. GURPP Framework
+2.1. Preliminaries
 Urban region.
@@
 The downstream prediction tasks aim to predict $y_{i}$ based on the region embedding $\zeta(r_{i})$ using a simple regression model like a Ridge Regression (Hoerl and Kennard, 1970).
-Figure 3.
-The Overview of our GURPP.
-(1) Create urban region graph that integrates specific urban entities.
-(2) Extract subgraph of each region according to the graph pattern.
-(3) Use multi-view learning to learn the subgraph representations with GURP model.
-(4) Adapt the region embeddings with manually-designed prompt method and task-learnable prompt learning method.
-3. Overview
-As shown in Figure 3, we propose a graph-based framework to pre-train and prompt urban region representations.
-Our approach begins with the creation of an urban region graph that integrates specific urban entities, providing a comprehensive structure for representing spatial and functional relationships within urban areas (discussed in Section 4.1).
-From this urban region graph, we extract subgraphs corresponding to each region, using graph patterns that characterize the region’s structure and interactions.
-These subgraph representations are then refined through multi-view learning using the GURP model, enabling the capture of generalizable information that can serve as the foundation for downstream tasks (explained in Section 4.2).
-To tailor these representations for specific tasks, we employ two prompt-tuning methods: a manually-designed prompt method and a task-learnable prompt learning method, each designed to incorporate task-specific knowledge into the pre-trained region embeddings (elaborated in Sections 5.1 and 5.2, respectively).
+However, learning effective region representations for specific tasks is challenging.
+First, task datasets are often very limited.
+This means the function $\zeta(\cdot)$ must work well in few-shot or even zero-shot situations.
+Second, urban spatial entities are diverse and large in number.
+This requires $\zeta(\cdot)$ to pull out general knowledge while still focusing on features specific to the task.
+2.2. Graph-based pre-training and prompting
+We propose GURPP, a unified framework that integrates urban graph pre-training with prompt learning.
+We formalize the GURPP framework in Eq.
+(1), where GURPP organizes spatial entities and inter-regional relationships within regions using a graph construction function $G(\cdot)$: During the pre-training stage $\mathcal{M}^{G}_{pre}$, the graph structure enables fine-grained feature extraction, such as capturing granular details like store brands within regions; In the prompting phase $\mathcal{M}^{G}_{pmt}$, GURPP dynamically assembles spatial entities and relationships for a region $r$ via graph prompts $\mathcal{P}$ to generate task-specific representations.
+(1) $\zeta(\mathcal{P},r;\mathcal{R})=\mathcal{M}^{G}_{pre}\left(\mathcal{M}^{G}_{pmt}\left(\mathcal{P},G(r)\right),G(\mathcal{R})\right)$
+GURPP integrates generalizable and task-specific knowledge within a unified framework.
+Its effectiveness can be demonstrated by recent advancements in graph prompt learning  (Wang et al., 2025; Yu et al., 2024b; Li et al., 2024b).
+Moreover, GURPP offers flexibility by supporting versatile forms of prompting for the same pre-trained model, such as manually-designed prompting (e.g., graph patterns  (Dong et al., 2017; Huang and Zitnik, 2020) and graph rules  (Fan et al., 2022a, b)) or task-learnable prompting (e.g.,  soft prompting  (Liu et al., 2024; Chen et al., 2023)), as long as the prompt operation conforms to the urban graph schema.
+In the following sections, we detail the GURPP framework with a four-step workflow in Figure 3.
+First, an urban graph is built to unify heterogeneous spatial entities (Section 3.1).
+Second, region-specific subgraphs are extracted using structural and interaction-based patterns to identify region-specific patterns and relationships within urban data (Section 3.2).
+Third, a pre-trained multi-view learning model refines these subgraphs into generalizable embeddings (Section 3.3).
+Finally, task-specific knowledge is injected via prompt-tuning strategies (Section 4), enabling efficient adaptation without compromising the pre-trained backbone.
 3. Pre-train General Representation
@@
 3.1. Urban Region Graph Construction
-An urban region graph is constructed to model the spatial structures and functional layouts of urban areas.
-The urban region graph, denoted as $G=(V,E,T_{V},T_{E})$, is a heterogeneous graph containing multiple types of nodes and relationships.
-Here, $v\in V$ and $e\in E$ represent a node and an edge, respectively, and $\phi(v)\in T_{V}$ $(\psi(e)\in T_{E}$) is a type function that assign a type to a specific node (edge).
-The urban region graph follows a schema that defines the relationships among various types of nodes and edges (see Figure 4).
+An urban region graph models the spatial structures and functional layouts of urban regions.
+Given a set of regions $\mathcal{R}$, the urban region graph, denoted as $G(\mathcal{R})=(V,E,T_{V},T_{E})$ (we simplify $G(\mathcal{R})$ to $G$ in the following), is a heterogeneous graph that captures multiple types of nodes and relationships.
+Here, $v\in V$ and $e\in E$ represent a node and an edge, respectively, while $\phi(v)\in T_{V}$ and $\psi(e)\in T_{E}$ are type functions that assign a specific type to a node or edge.
+The urban region graph follows a schema that defines the relationships among various types of nodes and edges (see in Appendix A.1).
 Without loss of generality, we consider three types of spatial entities: POIs, roads, and junctions.
-Each spatial entity is categorized: for instance, POI categories include shops and parks, road categories encompass highways and residential streets, and junction categories cover features like roundabouts.
-Figure 4. Schema of urban region graph.
+Each spatial entity carries rich semantics; for instance, POIs have categories like shops and parks, and brands such as Dunkin’ and Starbucks.
+Roads include types such as highways and streets and junctions cover categories like roundabouts.
+This design effectively captures spatial structures and functional layouts at a granular level.
+For instance, two nearby regions containing coffee shops with brands of Blue Bottle and Dunkin’ are treated the same in existing statistic-based methods due to identical POI categories.
+In contrast, our region graph leverages semantic relations like BrandOf to distinguish them, enabling fine-grain region understanding, such as identifying high-end vs.
+casual commercial areas.
 We construct the urban region graph from POI datasets and online maps.
-Since the urban region graph contains heterogeneous nodes and edges, we adopt the knowledge graph embedding technique TransR (Lin et al., 2015) to initialize the representations of all nodes and edges.
-This aligns the heterogeneous information into a unified vector space.
+Since the graph contains heterogeneous nodes and edges, we first adopt the knowledge graph embedding technique TransR (Lin et al., 2015) to initialize the representations of all nodes and edges, aligning the heterogeneous information into a unified vector space.
 3.2. Subgraph-centric Region Embedding
@@
 The region-induced subgraph is a subgraph of $G$ that includes the region node $r$ and matches the graph pattern $\mathcal{P}=(P_{V},P_{E},\theta)$.
-Here, $P_{V}\subseteq T_{V}$ and $P_{E}\subseteq T_{E}$ are sets of node and edge types, respectively, which are used to filter entities and relationships specifically useful for region embeddings.
+Here, $P_{V}\subseteq T_{V}$ and $P_{E}\subseteq T_{E}$ are sets of node and edge types, respectively, which filter entities and relationships specifically useful for region embeddings.
 The set $\theta\subseteq T_{V}$ indicates the node types that serve as termination points for subgraph extraction.
@@
 We extract $g_{r}$ recursively, starting with $V_{r}=\{r\}$ and $E_{r}=\varnothing$ and repeatedly expanding $V_{r}$ and $E_{r}$ until the termination points are reached.
-Specifically, we use a breadth-first search (BFS) starting from $r$ to expand the subgraph.
-During each iteration of BFS, we explore edges $e(u,v)$ where one endpoint $u$ is in $V_{r}$ and the other endpoint $v$ is not yet in $V_{r}$.
-We then check whether the type of $e$ is in $P_{E}$ and whether the type of the unseen node $v$ is not in the termination type set $\theta$.
-If $e$ passes the checks, we add it to the set $\Delta E_{r}$.
-Using $\Delta E_{r}$, we construct the node set $\Delta V_{r}$, which contains the endpoints in $\Delta E_{r}$ that are not already in $V_{r}$.
-Formally, the sets of incremental edges (i.e.,  $\Delta E_{r}$) and nodes (i.e.,  $\Delta V_{r}$) are as follows: $\Delta E_{r}=\{e(u,v)~{}|~{}u\in V_{r}\land\phi(v)\notin\theta\land e\in E\land\psi(e)\in P_{E}\}$ and $\Delta V_{r}=\{v|~{}v\notin V_{r}\land\exists u\in V_{r},e(u,v)\in\Delta E_{r}\}$.
-We then update the region-induced subgraph by setting $V_{r}=V_{r}\cup\Delta V_{r}$ and $E_{r}=E_{r}\cup\Delta E_{r}$.
-The expansion terminates if $\Delta E_{r}$ and $\Delta V_{r}$ are empty.
+Here, the maximum boundary for subgraph extraction is set to the two-hop neighbors of the region node according to the graph schema.
 3.2.2. Region subgraph encoding.
-In Section 4.1, each node in the subgraph $g_{r}$ has been assigned a global initial embedding using TransR.
-Here, we encode each subgraph $g_{r}$ into a vector $\vec{h}_{r}$ to further incorporate the local features of region $r$.
+In Section 3.1, each node in the subgraph $g_{r}$ has been assigned a global initial feature vector $\vec{v}$ using TransR.
+We then encode each subgraph $g_{r}$ into a vector $\vec{h}_{r}$ to further incorporate the local features of region $r$.
 Specifically, we use a heterogeneous graph encoder $\textsf{gEnc}(\cdot)$ to extract the heterogeneous and generic patterns within $g_{r}$, ultimately outputting the contextualized embedding of the subgraph.
-Our encoder $\textsf{gEnc}(\cdot)$ chooses the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) as the subgraph node encoder due to HGT’s capability of capturing node and edge-types semantics.
+We choose the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) as the subgraph node encoder $\textsf{gEnc}(\cdot)$ due to its capability of capturing node and edge-type semantics.
 By using HGT, $\textsf{gEnc}(\cdot)$ characterizes the heterogeneity of input graph by maintaining a set of node and edge-type dependent parameters, and the output embedding of each node $\vec{v}^{HGT}$ that integrates the global initial embeddings generated by TransR, and node and edge-type semantics of the subgraph $g_{r}$.
 This trait aligns with our need to capture the semantic features of heterogeneous entities and relationships within each region-induced subgraph.
-After obtaining the representations of each node in the subgraph, we further obtain the graph-level representation of $g_{r}$ through two steps of aggregation.
-For each subgraph $g_{r}$, we first use $\text{SUM}(\cdot)$ function to aggregate the representations of nodes of the same type to integrate the impact of the number of entities on the representation, obtaining the node-type embedding $\vec{\tau}_{i}$:
-$\vec{\tau}_{i}=\sum_{\phi(v)=\tau_{i}}\vec{v}^{HGT},\quad\tau_{i}\in T_{V},\forall v\in V_{r}.$
-Then we concatenate all representations of node types, and encode them with a Linear layer to capture the interaction between various types of nodes:
-$\vec{h}_{r}=\text{Linear}\left(\text{CONCAT}_{|T_{V}|}(\vec{\tau}_{i})\right).$
-The output $\vec{h}_{r}$ of $\textsf{gEnc}(\cdot)$ is the final embedding of region-induced subgraph $g_{r}$.
+Obtaining the representations of each node in the subgraph, we further derive the graph-level representation of $g_{r}$ through two steps of aggregation.
+For each subgraph $g_{r}$, we first use $\text{SUM}(\cdot)$ function to aggregate the representations of nodes of the same type to integrate the impact of the number of entities on the representation, obtaining the node-type embedding $\vec{\tau}_{i}$: $\vec{\tau}_{i}=\sum_{\phi(v)=\tau_{i}}\vec{v}^{HGT}$, $\tau_{i}\in T_{V},\forall v\in V_{r}$.
+Then we concatenate all representations of node types, and encode them with a Linear layer to capture the interaction between various types of nodes: $\vec{h}_{r}=\text{Linear}\left(\text{CONCAT}_{|T_{V}|}(\vec{\tau}_{i})\right).$ The output $\vec{h}_{r}$ of $\textsf{gEnc}(\cdot)$ is the final embedding of region-induced subgraph $g_{r}$.
 Once the graph encoder is pretrained, any graph that conforms to the pattern $\mathcal{P}$ can be encoded by $\textsf{gEnc}(\cdot)$ to generate the corresponding representations, demonstrating high flexibility.
-4.3. Multi-view Representation Learning
-We learn the region representation with multiple views to characterize general region features by using the spatial proximity, region satellite imagery and region-wise traffic information.
+3.3. Multi-view Self-supervised Learning
+We learn region representations in a self-supervised manner by integrating multiple views such as spatial proximity, region imagery and mobility patterns to characterize general region features.
 3.3.1. Spatial proximity.
@@
 We input $g_{r}$, $g_{r}^{+}$, and $g_{r}^{-}$ to the graph encoder $\textsf{gEnc}(\cdot)$, obtaining the vector representations $\vec{h}_{r}$, $\vec{h}_{r}^{+}$, and $\vec{h}_{r}^{-}\in\mathbb{R}^{d}$, respectively.
-Then we use triplet loss to minimize the distance between adjacent regions and maximize it between distant regions, with the loss function calculated:
+Then we use triplet loss to minimize the distance between adjacent regions and maximize it between distant regions, where $\delta$ represents the margin parameter of triplet loss:
 $\mathcal{L}_{sp}=\sum_{r\in\mathcal{R}}\max\{||\vec{h}_{r}-\vec{h}_{r}^{+}||_{2}-||\vec{h}_{r}-\vec{h}_{r}^{-}||_{2}+\delta,0\}.$
-where $\delta$ represents the margin parameter of triplet loss.
 3.3.2. Region imagery.
@@
 For a given region $r$, its imagery can be represented as the set $\{a_{1},a_{2},\cdots,a_{m}\}$.
-We employ a pre-trained visual model to extract visual features of the region $\vec{h}_{r}^{img}$ and fuse these features with the region’s representation with the contrastive loss $\mathcal{L}_{img}$ (see Appendix A.2 for more details).
-4.3.3. Human mobility.
-Human mobility describes the dynamics of urban systems and is typically represented by trajectory data.
+We employ a pre-trained visual model to extract visual features of the region $\vec{h}_{r}^{img}$ and fuse these features with the region’s representation with the contrastive loss $\mathcal{L}_{img}$ (see Appendix A.2 for details).
+3.3.3. Mobility pattern.
+Mobility pattern describes the dynamics of urban systems and is typically represented by trajectory data.
 For a given region $r$, the outflow feature $\mathbf{s}_{r}$ and inflow feature $\mathbf{t}_{r}$ are defined by calculating the number of outgoing and incoming trips within each time period.
 This results in the outflow and inflow feature sets $\{\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{N}\}$ and $\{\mathbf{t}_{1},\mathbf{t}_{2},\cdots,\mathbf{t}_{N}\}$, respectively, where $\mathbf{s}_{i},\mathbf{t}_{i}\in\mathbb{R}^{l}$, and $l$ is the number of intervals (e.g., 24).
-We then encode the inflow and outflow features using an encoder (such as a Multi-Layer Perceptron) to obtain $d$-dimensional embeddings:
-$\vec{h}^{\text{src}}_{r}=\textsf{Encoder}(\mathbf{s}_{r})\quad,\quad\vec{h}^{\text{dst}}_{r}=\textsf{Encoder}(\mathbf{t}_{r}).$
-Here we adopt the reconstruction loss $\mathcal{L}_{flow}$ to optimize the two embeddings (see Appendix A.3 for more details).
-4.3.4. Region representation learning.
+We encode the inflow and outflow features with an encoder (such as a Multi-Layer Perceptron) to obtain $d$-dimensional embeddings: $\vec{h}^{\text{src}}_{r}=\textsf{Encoder}(\mathbf{s}_{r})$ and $\vec{h}^{\text{dst}}_{r}=\textsf{Encoder}(\mathbf{t}_{r}).$ Here we adopt the reconstruction loss $\mathcal{L}_{flow}$ to optimize the two embeddings (see Appendix A.3 for details).
+3.3.4. Self-supervised region representation learning.
 We aim to jointly learn region representations from spatial proximity, imagery, and flow views to extract general urban region knowledge and embed this into a unified representation space.
@@
 $\mathcal{L}_{fuse}=\sum_{r\in\mathcal{R}}\sum_{k=1}^{4}\|\vec{h}^{k}_{r}-\textsf{Decoder}^{k}(\hat{h}_{r})\|_{2}^{2}.$
-Here, $\vec{h}^{k}_{r}$ ($k=1,2,3,4$) represents one of the four embeddings, and $\textsf{Decoder}^{k}$ is the corresponding decoder (such as a Multi-Layer Perceptron).
+Here, $\vec{h}^{k}_{r}$ ($k=1,2,3,4$) represents one of the four embeddings, and $\textsf{Decoder}^{k}$ is the corresponding decoder (implemented by a linear layer).
+This fusion module is easy to be extended to more views.
 The final loss function can be formulated as follows:
@@
 where $\mu$ is a parameter that balances the weight of the fusion loss.
-5. Prompt Task-specific Knowledge
-After obtaining a general region representation, a critical issue lies in how to integrate task-specific knowledge into pre-trained representations.
-To achieve this, we propose two graph-based prompt tuning methods: a manually-designed prompt method based on explicit domain knowledge and a task-learnable prompt learning method to extract implicit task-specific knowledge.
+4. Prompt with Task-specific Knowledge
+We propose two graph-based prompt tuning methods, a manually-designed prompt method and a task-learnable prompt learning method, to extract explicit and implicit task-specific knowledge.
 4.1. Manually-designed Prompt
-In the manually-designed prompt, we adjust the subgraph structure, i.e., , defining the graph pattern $\mathcal{P}(P_{V},P_{E},\theta)$ in Section 4.2, to emphasize specific relationships and entities relevant to the target task, as shown in Figure 5.
-We further enhance this process by manually adjusting the proportion of different types of edges and entities in the subgraph, thereby influencing the resultant embeddings.
+In the manually-designed prompt, we adjust the subgraph structure, i.e., defining the graph pattern (i.e., prompt) $\mathcal{P}(P_{V},P_{E},\theta)$ in Section 3.2, to emphasize specific relationships and entities relevant to the target task, as shown in Figure 4.
+Figure 4. Manually-designed prompt.
+We enhance this process by manually adjusting the proportion of different types of edges and entities in the subgraph, thereby influencing the resultant embeddings.
 To achieve this, we first define a set of task-specific entity weights $\mathbf{W}_{task}$ that reflect the importance of various types of entities for the downstream task.
-These weights are designed based on domain knowledge and the specific requirements of the task.
-Formally, we define $\mathbf{W}_{task}$ as:
-$\mathbf{W}_{task}=\{w_{\tau_{i}}\mid\tau_{i}\in P_{V}\}.$
-where $P_{V}$ is the set of all possible node types in the region-induced subgraph, and $w_{\tau_{i}}\in[0,1]$ represents the importance of entity type $\tau_{i}$.
-The weights $w_{\tau_{i}}$ are manually defined based on domain knowledge and are used as the deletion proportion for each entity type.
-We define a function $\textsf{Adjust}(\cdot)$ to delete nodes based on $\mathbf{W}_{task}$$\textsf{Adjust}(g_{r},\mathbf{W}_{task})\rightarrow g^{\prime}_{r}$ where $g_{r}=(V_{r},E_{r})$ is the original subgraph.
+These weights are designed based on the domain knowledge and specific requirements of task.
+Formally, we define $\mathbf{W}_{task}=\{w_{\tau_{i}}\mid\tau_{i}\in P_{V}\},$ where $P_{V}$ is the set of all possible node types in region-induced subgraph, and $w_{\tau_{i}}\in[0,1]$ represents the importance of entity type $\tau_{i}$.
+The weights $w_{\tau_{i}}$ are manually defined and used as the deletion proportion for each entity type.
+We define a function $\textsf{Adjust}(\cdot)$ to delete nodes based on $\mathbf{W}_{task}$$\textsf{Adjust}(g_{r},\mathbf{W}_{task})\rightarrow g^{\prime}_{r}$, where $g_{r}=(V_{r},E_{r})$ is the original subgraph.
+The adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ is input to the pre-trained GURP to generate the task-specific embeddings.
 The function Adjust obtains the adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ as follows:
@@
 (2) Edge Deletion: After node deletion, delete edges connected to the removed nodes, such that $E^{\prime}_{r}=\{e\mid e=(u,v)\in E;u,v\in V^{\prime}_{r}\}$.
-The adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ is then input to the pre-trained GURP to generate the task-specific embeddings.
-In summary, the manually-designed prompt is the graph pattern $\mathcal{P}$ with the set of task-specific entity weights $\mathbf{W}_{task}$, which guide the $\textsf{Adjust}(\cdot)$ function in modifying the subgraph $g_{r}$ to $g^{\prime}_{r}$.
-By incorporating these manually-designed prompts, we can tailor the urban region representations to better capture task-relevant entities and relationships.
-This approach also offers the advantage of adapting to downstream tasks without the need for additional training.
-It simply requires adjusting the input subgraph based on the manually-defined prompts.
-Figure 5. Manually-designed Prompt.
+The manually-designed prompt is the graph pattern $\mathcal{P}$ with the set of task-specific entity weights $\mathbf{W}_{task}$, which guide the $\textsf{Adjust}(\cdot)$ function in modifying the subgraph $g_{r}$ to $g^{\prime}_{r}$.
+Incorporating such prompts, we can tailor the urban region representations to better capture task-relevant entities and relationships.
 4.2. Task-learnable Prompt
-Apart from leveraging explicitly designed task-specific knowledge prompts to enrich region representations, in the presence of sufficient task data, we can also learn prompts based on the task data and pre-trained representations to induce implicit knowledge within the task data, thereby enhancing region representations.
-Considering that we describe the attributes of an urban region through subgraph induced with region nodes from $G$, it is intuitive to similarly describe the task adaptation to different region attributes with graphs as well.
-Therefore, we propose a prompt learning module to use prompt graphs to portray the adaptation of downstream task relative to the original region embedding, as shown in Figure 6.
+Apart from leveraging explicitly designed task-specific knowledge prompts to enrich region representations, we can also learn prompts based on the task data and pre-trained representations to induce implicit knowledge within the task data, thereby enhancing region representations.
+Considering that we describe the attributes of an urban region through subgraph induced with region nodes from $G$, it is intuitive to describe the task adaptation to different region attributes with graphs as well.
+Therefore, we propose a prompting module to use prompt graphs to portray the adaptation of downstream task relative to the original region embedding, as shown in Figure 5.
 Figure 5. Task-learnable prompt.
-We introduce the graph kernel method (Vishwanathan et al., 2010; Nikolentzos and Vazirgiannis, 2020; Feng et al., 2022) to assess the similarity of between the prompt graphs and the region subgraphs due to its excellent performance and computational efficiency, then enhancing the pre-trained region embedding.
-Let $\mathcal{G}$ denote the prompt graph set consisting of $m$ trainable (hidden) prompt graphs, where each prompt graph $G^{p}\in\mathcal{G}$ is defined as $G^{p}=(V^{p},E^{p},\mathbf{X}^{p})$.
-Here, $\mathbf{X}^{p}\in\mathbb{R}^{|V^{p}|\times d}$ represents the node attributes, with $d$ being the dimension of these attributes.
-For each region subgraph $g_{r}$, we utilize an attribute graph kernel function $K_{p}(\cdot,\cdot)$ to measure the similarity between $g_{r}$ and each prompt graph $G^{p}_{1},G^{p}_{2},\ldots,G^{p}_{m}$.
-Specifically, we employ the $P$-step random walk kernel (more details in Appendix A.4) to compute the kernel value between the region subgraph $g_{r}$ and each prompt graph $G^{p}_{i}$.
-These values are concatenated to form an $m$-dimensional prompt vector $\vec{h}_{p}\in\mathbb{R}^{m}$ for each region.
-We concatenate the prompt vector $\vec{h}_{p}\in\mathbb{R}^{m}$ before the pre-trained region representation $\vec{h}_{r}$ to derive region embedding $\vec{h}^{\prime}_{r}=\vec{h}_{p}||\vec{h}_{r}$.
-The new embeddings are then fed into a linear layer $\textsf{Linear}(\cdot)$.
-We update the prompt graphs and linear layer parameters, by optimizing on the downstream labeled task, to obtain the final prompt representation with the simple MSE loss function:
-$\mathcal{L}_{task}=\frac{1}{N}\sum_{i=1}^{N}(y^{pred}_{i}-y^{label}_{i})^{2}.$
+We introduce the graph kernel method (Vishwanathan et al., 2010; Feng et al., 2022) to measure the similarity between the prompt graphs and the region subgraphs due to its performance and computational efficiency.
+Let $\mathcal{P}$ denote the set (i.e., prompt) consisting of $m$ trainable (hidden) prompt graphs, where each graph $p\in\mathcal{P}$ is defined as $p=(V^{p},E^{p},\mathbf{X}^{p})$.
+Here, $\mathbf{X}^{p}\in\mathbb{R}^{|V^{p}|\times d}$ denotes the node attribute matrix, where $d$ is the attribute dimension.
+For each region subgraph $g_{r}$, we utilize an attribute graph kernel function $K_{p}(\cdot,\cdot)$ to measure the similarity between $g_{r}$ and each prompt graph $p_{1},p_{2},\ldots,p_{m}$.
+Specifically, we employ the $P$-step random walk kernel (more details in Appendix A.4) to compute the kernel value between region subgraph $g_{r}$ and each prompt graph $p_{i}$.
+These values are concatenated to form an $m$-dimensional prompt vector $\vec{h}_{\mathcal{P}}\in\mathbb{R}^{m}$ for each region.
+The prompt vector is used as a data operator to modify the pre-trained graph model.
+Conceptually, we can concatenate $\vec{h}_{\mathcal{P}}$ with each initial node feature vector $\vec{v}$ and feed the result into a linear layer $\textsf{Linear}(\cdot)$, such that each entity is modified by $\vec{v}_{*}=\textsf{Linear}(\vec{h}_{\mathcal{P}}||\vec{v})$.
+To reduce the complexity, in practice, we concatenate $\vec{h}_{\mathcal{P}}$ with the pre-trained region embedding $\vec{h}_{r}$ instead of each entity embedding.
+To learn the prompting model, we update the prompt graphs and linear layer based on specific downstream task with a simple MSE loss function.
+Table 1.
+Main performance comparison of different models on downstream tasks.
+The best results for each metric are highlighted in bold, the second-best are underlined, and the third-best are marked with an asterisk.
+Model New York City (NYC) Chicago (CHI)
+Crime Check-in Carbon Income Crime Crash Carbon Education
+MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
+TransR-N 111.32 148.86 534.70 805.28 115.65 175.73 36799 45098 209.07 307.45 106.91 176.86 97.15 126.59 653.68 932.56
+TransR-G 109.22 146.91 530.43 802.43 116.80 179.25 36363 44571 208.40 306.16 104.68 174.42 95.76* 124.37* 601.97 875.24
+node2vec 138.46 175.80 573.12 837.44 174.45 225.42 45330 56182 246.07 349.97 121.90 174.91 115.62 148.95 832.89 1139.33
+GAE 96.85 126.36 349.94 542.92 115.99 174.35 35338 43200 200.61* 300.24 100.54 173.32 95.29 123.66 702.82 986.70
+MVURE 72.47* 99.84 310.97 498.08 99.30 151.01 24031 35351 206.08 289.86* 96.95 148.87* 98.19 130.31 593.80 843.10
+MGFN 73.50 97.44* 300.15 456.63* 136.49 185.82 30266 40717 214.02 310.83 108.37 179.85 96.56 125.50 709.91 998.03
+HRE 76.93 102.07 329.55 513.80 102.95 160.58 25903 35418 206.61 298.39 99.95 158.63 97.11 133.00 555.56 806.60
+HREP 74.85 100.04 320.34 505.52 102.49 159.68 24643 34148 206.59 298.24 99.77 158.66 97.19 132.91 555.01 806.36
+ReCP 73.85 100.56 292.87* 524.89 95.55 150.84 24440* 35044 207.49 298.97 93.30* 157.52 95.84 125.73 455.98* 708.37*
+GURP 72.09 93.22 271.69 412.92 108.17 158.03 25268 34943* 161.27 225.24 80.36 108.78 97.83 133.03 455.43 677.36
+GURPP 71.48 91.61 261.45 388.23 100.53* 152.70* 22902 31945 146.39 203.05 65.30 92.11 64.844 88.56 420.62 656.98
+Enhance 1.36% 5.98% 10.73% 14.98% -4.95% -1.21% 4.70% 6.45% 27.03% 29.95% 30.01% 38.12% 31.95% 28.39% 7.76% 7.26%
 5. Experiments
-We conduct extensive experiments to investigate the following research questions (RQs):
-  • RQ1: Can our proposed pre-train and prompt framework (GURPP) outperform previous approaches?
-  • RQ2: Does GURP have the ability to capture and apply heterogeneous and generic patterns?
-  • RQ3: How does the manually-designed prompt work?
-  • RQ4: How does the task-learnable prompt perform and how does it look like?
-  • RQ5: How does each component contribute to GURPP?
+We conduct extensive experiments on six downstream tasks of two cities from the following perspectives: (1) overall performance comparison with SOTA baselines, (2) effectiveness analysis of two prompting methods, (3) few-shot and zero-shot generalization capabilities, (4) ablation studies, (5) scalability evaluation, and (6) comparison with language-based prompting methods.
 5.1. Experimental Setup
 5.1.1. Datasets.
-We collect the datasets of region division, POI, urban imagery, taxi trips, road network, check-in, crime and crash for two representative cities in the United States: New York City (NYC) and Chicago (CHI).
-The detailed statistics and sources of the dataset is available in Appendix A.5.
+We collect the datasets of region division, POI, urban imagery, taxi trips, road network and six downstream tasks for two representative cities in the United States: New York City (NYC) and Chicago (CHI).
 5.1.2. Baselines.
-We compare our model with the following representative baseline models: knowledge graph embedding methods and its variants (TransR (Lin et al., 2015), TransR-N, TransR-G), graph embedding methods (node2vec (Grover and Leskovec, 2016), GAE (Kipf and Welling, 2016)), and urban region embedding methods (MVURE (Zhang et al., 2021), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023a), HRE, ReCP (Li et al., 2024)).
-These baselines cover the range from traditional graph embedding to complex models that take into account urban-specific attributes, providing us with a comprehensive performance comparison benchmark.
-Specific details and descriptions can be found in the Appendix A.6.
+We compare our model with the following representative baselines: knowledge graph embedding methods and its variants (TransR (Lin et al., 2015), TransR-N, TransR-G), graph embedding methods (node2vec (Grover and Leskovec, 2016), GAE (Kipf and Welling, 2016)), and urban region embedding methods (MVURE (Zhang et al., 2021), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023a),ReCP (Li et al., 2024a)).
+We follow the descriptions in these works to preprocess the data into the corresponding format and run the baselines with the parameter settings specified in the papers to ensure fairness.
 5.1.3. Metrics and Implementation.
-To evaluate the prediction performance, we use three widely recognized evaluation metrics: the coefficient of determination ($R^{2}$), root mean squared error (RMSE), and mean absolute error (MAE).
+We use three widely recognized evaluation metrics to evaluate the prediction performance: the coefficient of determination ($R^{2}$), root mean squared error (RMSE), and mean absolute error (MAE).
 Better performance is indicated by a higher $R^{2}$ and lower RMSE and MAE values.
-Our model includes three version: 1) GURP represents the pre-train module, 2) $\text{GURPP}_{M}$ is GURP with manually-designed prompt, and 3) $\text{GURPP}_{T}$ is GURP with task-learnable prompt.
-For detailed implementation, please refer to Appendix A.7 and our provided code repository.
-Table 1.
-Main performance comparison of different models on downstream tasks across NYC and CHI datasets.
-The best results for each metric are highlighted in bold, the second-best are underlined, and the third-best are marked with an asterisk.
-The last and the third-to-last rows show the percentage improvement of the GURPP and GURP models over the third-best model.
-Model New York City (NYC) Chicago (CHI)
-Crime Prediction Check-in Prediction Crime Prediction Crash Prediction
-MAE RSME $R^{2}$ MAE RSME $R^{2}$ MAE RSME $R^{2}$ MAE RSME $R^{2}$
-TransR-N 111.319 148.863 -0.014 534.699 805.282 0.031 209.072 307.454 -0.037 106.914 176.860 0.001
-TransR-G 109.221 146.905 0.013 530.427 802.426 0.038 208.404 306.155 -0.028 104.680 174.419 0.028
-node2vec 138.458 175.797 -0.414 573.118 837.441 -0.048 246.066 349.965 -0.343 121.900 174.907 0.023
-GAE 96.846 126.364 0.270 349.942 542.920 0.560 200.606* 300.237 0.011 100.536 173.323 0.040
-MVURE 72.471* 99.844 0.544 310.973 498.076 0.629 206.080 289.859* 0.079* 96.953 148.865* 0.292*
-MGFN 73.502 97.440* 0.566* 300.150 456.626* 0.688* 214.021 310.832 -0.059 108.3740 179.854 -0.033
-HRE 76.926 102.07 0.523 329.547 513.801 0.606 206.612 298.389 0.024 99.952 158.625 0.196
-HREP 74.854 100.035 0.542 320.335 505.516 0.618 206.591 298.238 0.025 99.766 158.658 0.196
-ReCP 73.853 100.558 0.537 292.871* 524.890 0.588 207.485 298.965 0.020 93.298* 157.515 0.207
-GURP 72.094 93.216 0.603 271.689 412.918 0.745 161.266 225.244 0.444 80.362 108.781 0.622
-Enhance 0.52% 4.33% 6.54% 7.23% 9.57% 8.28% 19.61% 22.29% 462.03% 13.87% 26.93% 113.01%
-GURPP 71.481 91.612 0.616 261.450 388.225 0.775 146.385 203.050 0.548 65.297 92.114 0.729
-Enhance 1.36% 5.98% 8.83% 10.73% 14.98% 12.65% 27.03% 29.95% 593.67% 30.01% 38.12% 149.66%
-6.2. RQ1 & RQ2: Overall Performance
-We evaluate the overall performance of baselines and our model on three downstream tasks.
-The results of the experiments are shown in Table 1, where the GURPP is all implemented by $\text{GURPP}_{T}$.
-From the results, we find the following:
-(1) GURPP demonstrates superior performance to baselines in both pre-training and prompting phases.
-In the pre-training phase, the performance of GURP improves by 6.54%, 8.28%, 462.03%, and 113.01% over the best baseline of $R^{2}$ on two urban datasets.
-Meanwhile, on all downstream tasks, the average performance on MAE and RMSE metrics improves by 17.28% and 22.26%, respectively.
-Subsequently, the performance of GURPP is further improved by an average of 9.50%, 9.28%, and 11.703% of MAE, RMSE and $R^{2}$, respectively, when we introduce the task-learnable prompt.
-The results validate the significant advantages of our proposed pre-train and prompt framework.
-(2) To validate the ability of our pre-training method to capture generic urban knowledge, we compare the performance of state-of-the-art urban region embedding methods on datasets from two distinct cities, NYC and CHI.
-The results show that while these baselines perform well on the NYC dataset, their performance drops significantly on the CHI dataset.
-For instance, the performance on CHI decreases by an average of 71.8% in $R^{2}$ compared to NYC.
-In contrast, our GURPP model demonstrates remarkable consistency, with only an 8.49% decrease in $R^{2}$.
+Our model includes three version: 1) GURP represents the pre-train module, 2) $\textbf{GURPP}_{M}$ is GURP with manually-designed prompt, and 3) $\textbf{GURPP}_{T}$ is GURP with task-learnable prompt.
+Details on datasets and implementation are in Appendix A.5.
+5.2. Overall Performance
+We evaluate the overall performance of the baselines and our model on datasets collected from two cities, covering six different downstream tasks.
+The results of the experiments are shown in Table 1, from which we find the following.
+(1) GURPP demonstrates superior performance across both cities and tasks.
+In the pre-training phase, GURP achieves state-of-the-art performance on 6 out of 8 tasks, with particularly significant improvements on CHI datasets (27.03%-31.95% MAE reduction on Crime/Crash prediction).
+When enhanced with prompting methods, GURPP further improves performance by an average of 11.30% (MAE) and 10.16% (RMSE) across all tasks compared to GURP.
+The consistent improvements across heterogeneous tasks (12.46% average MAE gain and 14.62% RMSE gain over baselines) validate the effectiveness of our pre-training and prompting framework.
+(2) The results highlight our model’s exceptional adaptability to different urban contexts.
+Specifically, GURPP achieves average improvements of 24.2% (MAE) and 25.9% (RMSE) in CHI versus 3.0% (MAE) and 6.6% (RMSE) in NYC.
 This may be due to the fact that the baselines tend to be effective only at extracting information specific to certain data types, but they do not consider spatial structure and functional layouts among urban entities.
-For example, existing baselines rely on the flow view to extract information.
-However, the content of flow view in the CHI dataset is sparse, leading to a significant degradation in the performance of downstream task.
-While GURPP shows excellent performance on both urban datasets, which demonstrates the robustness and universality of GURPP across diverse urban environments.
-(3) To further validate the effectiveness of our subgraph-centric and multi-view fusion module, we conduct experiments comparing the performance of models with and without subgraphs and multi-view fusion.
-Specifically, we compare the TransR-G model (adopts region-induced subgraph and regards graph-level embedding as the region embedding) with the TransR-N model (directly uses region nodes embedding as region embedding).
-The results indicate that TransR-G outperforms TransR-N by an average of 1.27% in MAE, underscoring the effectiveness of the subgraph-centered approach.
-This improvement is attributed to the subgraph-centric method’s ability to better preserve the structural and contextual information of regions, providing more accurate and meaningful urban region embeddings.
-The average 32.55% improvement in MAE for GURP over TransR-G is mainly due to the multi-view fusion strategy, which improves the model’s comprehensive understanding of urban features.
-6.3. RQ3: Manually-designed Prompt Study
-In the real world, the incidence of crime is often closely linked to specific POI categories, roads and junctions, for example, ATMs, bars, parks, etc.
-are more likely to be the location of a crime.
-The density of roads and the complexity of junctions are also important factors affecting crime rates.
-As for the check-in behavior, the brand of the POI is given more attention, and certain users may be more inclined to check in at locations of well-known brands such as Starbucks, McDonald’s, or other specific brands.
-Therefore, we designed two patterns as shown in Figure 7 to validate the performance of manually-designed prompt.
-The results are shown in Table 2, from which it can be found that crime prediction does pay more attention to factors such as POI categories, roads and junctions, which are closely related to criminal behavior.
-Meanwhile, removing POI brand information shows a significant decrease in the performance of check-in prediction.
-These results also demonstrate that the prediction performance of the model can be improved by manually-designed prompt.
-Figure 7.
-The structure of the manually-designed prompt based on Pattern 1 (P1) and Pattern 2 (P2).
-P1 ignores POI brand information.
-P2 randomly samples POI, road, and junction nodes around the region with a weight of 0.9.
-Table 2. Performance of $\text{GURPP}_{M}$.
+For example, existing baselines rely on the flow view for information extraction, but the sparse content in the CHI dataset significantly degrades downstream task performance.
+In contrast, GURPP performs excellently on both urban datasets, demonstrating its robustness and universality across diverse urban environments.
+(3) To further validate the effectiveness of our subgraph-centric and multi-view fusion module, we conduct experiments comparing the performance of models with and without these modules.
+Specifically, the result that TransR-G outperforms TransR-N by an average of 1.27% (MAE) indicates that the subgraph-centric method can preserve the structural and contextual information of regions, providing more effective urban region embeddings.
+The average 32.55% (MAE) improvement for GURP over TransR-G is mainly due to the heterogeneous feature extraction, which improves the model’s comprehensive understanding of urban features.
+5.3. Effectivness of Prompting Methods
+5.3.1. Manually-designed Prompt Study.
+In the real world, crime rates are closely linked to specific POI categories, roads, and junctions, with locations like ATMs, bars, and parks being more crime-prone.
+Road density and junction complexity also impact crime.
+Regarding user check-ins, brand popularity plays a key role, with users often preferring well-known brands like Starbucks or McDonald’s.
+Based on these observations, we design two patterns: P1 (ignoring POI brand information) and P2 (randomly sampling POI, road, and junction nodes with a 0.9 ratio) to evaluate the performance of manually-designed prompts across different tasks.
+Table 2. Performance of manually-designed prompt $\text{GURPP}_{M}$.
 Model Crime Prediction (NYC) Check-in Prediction (NYC)
-MAE RSME $R^{2}$ MAE RSME $R^{2}$
+MAE RMSE $R^{2}$ MAE RMSE $R^{2}$
 GURP 72.094 93.216 0.603 271.689 412.918 0.745
@@
 $\text{GURPP}_{M}$-P2 69.421 91.834 0.614 262.437 397.466 0.764
-6.4. RQ4: Task-learnable Prompt Study
-6.4.1. Effectiveness of Task-learnable Prompt.
-We utilize GURP pre-trained embedding with the simple prefix prompt (proposed by HREP(Zhou et al., 2023a)) and task-learnable prompt respectively for comparative analysis.
-Both of these prompt methods are data-driven and seek to improve the region representation by learning data for specific downstream tasks.
+Results in Table 2 show that focusing on factors such as POI categories, roads, and junctions, closely related to criminal behavior, can improve crime prediction performance.
+Meanwhile, removing POI brand information shows a significant decrease in the performance of check-in prediction.
+This demonstrates that our manually-designed prompting method can improve the prediction performance of downstream tasks with explicit knowledge injection.
+5.3.2. Task-learnable Prompt Study.
+We utilize GURP pre-trained embedding with the simple prefix prompt (proposed by HREP(Zhou et al., 2023a) denoted as GURP+Pre), which directly concatenates hidden vector before pretrained embedding without modeling task-adaptation, and our task-learnable prompt for comparative analysis.
+Both of these prompt methods are data-driven and seek to improve the region representation for specific downstream tasks.
 We conduct experiments on crime prediction task in NYC and crash prediction task in CHI, respectively, and the results are shown in Table 3.
-Our task-learnable prompt outperforms pre-trained embeddings across both tasks, with improvements of 8.83% of $R^{2}$ in crime prediction and 149.66% of $R^{2}$ in crash prediction, indicating that the prompt is highly effective in providing tailored guidelines for different downstream tasks.
-However, in crime prediction task, the prefix prompt have an negative effect, resulting in a performance decrease of 1.16% of $R^{2}$.
-This demonstrates that our task-learnable prompt is more stable and effective compared to prefix prompt.
-Table 3. Effectiveness of Task-learnable Prompt.
+Results show that our task-learnable prompt outperforms pre-trained embeddings across both tasks, with improvements of 2.16% ($R^{2}$) in crime prediction and 17.20% ($R^{2}$) in crash prediction, indicating that the prompt is highly effective in providing tailored guidelines for different downstream tasks.
+In addition, $\text{GURPP}_{T}$ outperforms the prefix prompt in both tasks.
+This is because the prefix prompt only utilizes the pre-trained region representations for prompting, while our graph-based prompting method also incorporates the original region subgraphs to prompt, better capturing the correlations between task and region attributes.
+Table 3. Performance of task-learnable prompt $\text{GURPP}_{T}$.
 Framework Crime Prediction (NYC) Crash Prediction (CHI)
-MAE RSME $R^{2}$ MAE RSME $R^{2}$
+MAE RMSE $R^{2}$ MAE RMSE $R^{2}$
 GURP 72.094 93.216 0.603 80.362 108.781 0.622
@@
 $\text{GURPP}_{T}$ 71.481 91.612 0.616 65.297 92.114 0.729
-6.4.2. Visualization of Learned Prompt Graph.
-Compared to prefix prompt, our task-learnable prompt is able to generate a series of prompt graphs during training, which enhances the clarity of our prompt method.
-We visualize prompt graphs that reveal key structural features in the urban region graph of interest in the task.
-In Figure 8, we show some prompt graphs for crime prediction tasks in CHI.
-(Additional prompt graphs can be found in Appendix A.8.) These results show that the prompt graphs generally exhibits circular structural features, implying that the model have identified and emphasized interactions between urban entities.
-Figure 8. Visualization of learned prompt graphs.
-6.5. RQ5: Ablation Studies
-Finally, we conduct the ablation studies in NYC dataset to investigate the effectiveness of the individual components of GURPP.
-Variants of GURPP are designed as follows: GURPP/P.
-We use only the pre-training process to learn urban region general representation.GURPP/F,  GURPP/S,  GURPP/I.
-We remove the flow view, spatial proximity view, and imagery view from GURPP, respectively.
-GURPP/K.
-We replace the knowledge graph embedding initialization with random initialization.
-Figure 9 presents the ablation results of our model and its variants.
+5.3.3. Comparison of Two Prompting Methods.
+We compare the performance of two proposed prompting methods across different cities and tasks to evaluate their effectiveness in various scenarios.
+In addition to the two prompting patterns P1 and P2, we also define two additional patterns: P3 (randomly samples POI with a weight of 0.9 and ignores road and junction category) and P4 (randomly samples POI with a weight of 0.7 and ignores road and junction), to further validate the effectiveness of different prompting methods.
+Results show in Table 4.
+Table 4. Comparison of two prompting methods.
+Method New York City (NYC) Chicago (CHI)
+Check-in Carbon Income Crime Carbon Education
+RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE
+GURP 412.92 0.23 25268 225.24 -0.33 455.43
+$\text{GURPP}_{T}$ 388.23 0.17 24745 203.05 0.41 420.62
+$\text{GURPP}_{M}$-P1 429.56 -0.13 27839 234.26 -0.35 465.81
+$\text{GURPP}_{M}$-P2 397.47 -0.22 26392 234.74 -0.40 458.59
+$\text{GURPP}_{M}$-P3 464.90 0.28 28170 256.53 -0.66 486.19
+$\text{GURPP}_{M}$-P4 340.25 0.16 22902 238.36 -0.49 502.86
+From the results, we observe that when the pre-trained city region data is relatively sparse (e.g., in cases of CHI), task-learnable prompting methods, especially our $\text{GURPP}_{T}$, achieve better enhancement effects.
+Conversely, when the data is more abundant (e.g., in cases of NYC), manually-designed prompts with explicitly injected knowledge perform better.
+Moreover, different manually-designed prompts explicitly capture distinct types of knowledge, leading to varying performance across different tasks, which demonstrates the flexibility of our proposed graph-based prompting approach.
+5.4. Few-shot & Zero-shot Performance
+We compare the performance of $\text{GURPP}_{T}$ and baselines in few-shot and zero-shot scenarios to validate the generalizability of pre-trained representations and superiority of prompting methods.
+5.4.1. Few-shot Prediction.
+With the pre-trained region representations as input, we randomly sample 10% of the region data as the training set to develop the prediction model, while the remaining 90% is designated as the testing set, to evaluate the performance of $\text{GURPP}_{T}$ with other baselines in few-shot scenarios.
+Figure 6(a) shows that $\text{GURPP}_{T}$ outperforms all baselines in the few-shot scenario.
+This is partly because the pre-trained representations extract generic regional features that exhibit better adaptability in low-resource settings, which, when combined with task-learnable prompts, can better accommodate few-shot prediction.
+Additionally, our prompting method shows effective transferability compared to HREP.
+5.4.2. Zero-shot Prediction.
+In this experiment, we first trained the prediction model on NYC crime datasets with NYC pre-trained region embeddings, and then evaluated it on CHI crime datasets with CHI pre-trained region embeddings without any fine-tuning.
+This setup enables a comparative analysis of the zero-shot transferability between our task-learnable prompt and various baselines.
+Results show in Figure 6(b).
+(a) Few-shot prediction.
+(b) Zero-shot prediction.
+Figure 6. Few-shot & zero-shot performance.
+Compared with baselines without any prompting methods (i.e., HRE, MVURE, and ReCP), models with prompting generally exhibit better zero-shot transferability, suggesting that prompting model can capture task-specific features independent of the city to some extent.
+Under the same prompting method (prefix prompt), GURP+Pre demonstrates better transferability than HREP, highlighting the superior generalizability of our GURP pre-trained representations.
+A comparison between GURPP and GURP+Pre reveals that our task-learnable prompting model outperforms the prefix prompt, which can be attributed to the handling of the original region subgraphs within our prompting model.
+5.5. Ablation Studies
+We perform ablation studies on the NYC dataset to evaluate the effectiveness of each individual component of our proposed model.
+GURPP includes both pretraining and task-learnable prompting module.
+GURP contains only the pretraining module.
+/F, /S, /I, /M denote the removal of the flow view, spatial proximity view, imagery view, and multi-view fusion from GURP, respectively.
+/K indicates replacing the knowledge graph embedding initialization with random initialization.
+Figure 7 presents the results.
 Spatial proximity view and knowledge graph embedding method significantly enhance the model’s performance, lifting it by 151.64% and 156.67% of $R^{2}$, respectively.
-These components effectively capture the heterogeneous and generic patterns of interactions among entities by elucidating the spatial relationships and structured knowledge of the urban region.
+These components capture the heterogeneous and generic patterns of interactions among entities by elucidating the spatial relationships and semantic knowledge of urban regions.
+The removal of the multi-view fusion module results in learned representations exerting varying effects across different tasks, which indicates that this module contributes to learning more stable representations.
 The flow view and the imagery view can also greatly enhance the performance of the model, increasing it by 29.04% and 21.11%, respectively.
@@
 Prompt learning, with an improvement of 3.12%, effectively elevates model performance by introducing guiding prompts that allow the model to delve deeper into understanding and utilizing task-relevant knowledge.
-Figure 9. Results of ablation studies.
+Figure 7. Ablation studies.
+5.6. Scalability Analysis
+We also validate the scalability of our graph-based method by measuring the pre-training and prompting tuning duration across various scales of regions and region subgraph nodes.
+Specifically, we use 180 regions in NYC as the base and evaluate the training duration of GURPP over 500 epochs and prompt tuning over 200 epochs for region scale factors of 0.1, 0.2, up to 5 times(1.6 times for pre-training), and evaluate the pretraining duration using 0.05, 0.1, up to 1.0 times the total number of nodes in each region subgraph.
+For prompt tuning, we conduct evaluations with subgraphs containing 1, 5, up to 120 nodes with the base number 50, which is also the actual experiments settings.
+The results are averaged over each epoch, shown in Figure 8.
+As shown in Figure 8(a), when the region subgraph size is fixed (GURP uses all nodes and prompt tuning uses 50 nodes per subgraph), the training time for pretraining exhibits a superlinear growth with the number of regions.
+This is mainly due to the flow feature extraction, where the model incorporates dynamics between all region pairs, whereas prompt tuning scales linearly with the number of regions.
+When the number of regions is fixed, the pretraining duration shows a sublinear relationship with the number of nodes in the regional subgraph, ensuring the scalability of GURP.
+Due to the random walk kernel, the training time for the task-learnable prompt scales superlinearly with the number of nodes.
+We set the number of nodes to 50, maintaining acceptable training efficiency.
+(a) Pre-training Time.
+(b) Prompt Tuning Time.
+Figure 8. Impact of data scale on training time.
+5.7. Case Study
+To validate the superiority of GURPP, we conduct a case study to predict the top 10 regions with the highest crime rates in NYC.
+The study compares our manually designed prompt, $\text{GURPP}_{M}$-P4, against large the language models (LLMs), including Deepseek R1 (www.deepseek.com) and Gemini 2 Flash (deepmind.
+google/technologies/gemini/flash/).
+Detailed descriptions are in Appendix A.5.4.
+As shown in the Figure 9, GURPP is able to predict crime regions (the regions in the upper half of the figure) but LLMs fail.
+This indicates that, compared to language prompts, which directly describe the task to induce relevant knowledge in large language models, GURPP explicitly incorporates task-specific knowledge into the representation using graphs, leading to better performance.
+Figure 9. Comparison with language prompt. Shaded regions represent potential crime regions.
 6. Related Work
-Extensive studies have been conducted to learn urban region representation from different perspectives. We combine the existing work categories in Table 4.
-Table 4. Cats.of Related Works on Urban Region Embedding
-Methods Task-specific learning or fine-tuning Prompting
-Non-graph (Yang et al., 2022; Zhou et al., 2023b; Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Jenkins et al., 2019) (Yuan et al., 2024; Yan et al., 2024; Zhou et al., 2023a)
-Graph (Xu et al., 2023; Zhang et al., 2024; Bai et al., 2019; Han et al., 2020; Zhang et al., 2021; Wu et al., 2022; Liu et al., 2023a; Ning et al., 2024) GURPP(Ours)
-Task-specific learning or fine-tuning.
+Urban Region Embedding.
 Urban region representation learning aims to extract complex patterns and relationships within urban regions and embed them into a representation space.
-Existing research can be broadly classified into two categories: task-specific supervised region representation learning (Bai et al., 2019; Han et al., 2020; Yang et al., 2022; Zhou et al., 2023b) and general unsupervised region representation learning (Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021).
-These works utilize feature aggregation models to generate representation from predefined region statistics with limitations in capturing detailed spatial layout and semantics.
-Moreover, using predefined statistical features makes it challenging to further adapt the representation to the specificities of various downstream tasks.
-We address the limitations by leveraging a urban region graph and subgraph-centric region embeddings, which learn the generic knowledge of urban spatial layout and generate flexible region representation tailored to different tasks.
+Existing research can be broadly classified into two categories: task-specific supervised region representation learning (Bai et al., 2019; Yang et al., 2022; Zhou et al., 2023b) and general unsupervised region representation learning (Li et al., 2023; Zhang et al., 2023; Li et al., 2024a; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021; Sun et al., 2024a; Xu and Zhou, 2024).
+These works, focusing more on exploring multi-view fusion methods to generate representation based on predefined region statistics features with limitations in capturing detailed spatial layout and semantics.
+Moreover, the lack of fine-grained characterization of region features makes it challenging to further adapt the representation to the specificities of downstream tasks.
+We address the limitations by using an urban region graph and subgraph-centric region embeddings.
+Although Zhao et al.
+(Zhao et al., 2023) employ a graph-based approach to model urban regions at a fine-grained level, their framework is not specifically designed for urban-region pre-training and prompting tasks.
 Pre-training and Prompting.
-Recent years have witnessed the significant success of pre-training and prompting paradigm in fields of NLP (Liu et al., 2023b; Brown et al., 2020; Roberts et al., 2020; Gao et al., 2021), CV (Yoo et al., 2023; Radford et al., 2021; Zhou et al., 2022b, a) and graph learning (Sun et al., 2023; Yu et al., 2024b; Ma et al., 2024; Yu et al., 2024a).
-Many works in urban computing have also made attempts in this regard (Xu et al., 2023; Zhang et al., 2024; Balsebre et al., 2023; Yuan et al., 2024; Yan et al., 2024).
+The Pre-training and prompting paradigms have achieved significant success in NLP (Liu et al., 2023c; Roberts et al., 2020), CV (Yoo et al., 2023; Zhou et al., 2022b, a), and graph learning (Sun et al., 2023; Yu et al., 2024c; Ma et al., 2024; Yu et al., 2024a; Liu et al., 2023b; Yang et al., 2024; Tang et al., 2024).
+Many works in urban computing have also made attempts in this regard (Xu et al., 2023; Zhang et al., 2024; Balsebre et al., 2023; Yuan et al., 2024; Yan et al., 2024; Xiao et al., 2024; Li et al., 2024c).
 For instance, Yuan et al.
@@
 Yan et al.
-(Yan et al., 2024) introduce large language models (LLMs) to describe urban imagery features and unify different prediction tasks through textual prompts.
-However, the urban tasks explored in these work differ from ours, with a distinct focus.
-Our work can also benefit from the integration of LLMs to enhance fine-grained characterization capabilities.
-As a similar work, HREP (Zhou et al., 2023a) introduces prefix prompt to region embedding, but the pre-trained model fails to adequately capture the fine-grained features of regions.
-We propose both manually-designed and task-learnable prompts built on graph-based pre-train model to address this issue.
+(Yan et al., 2024) introduce LLMs to describe urban imagery features and unify different prediction tasks through textual prompts.
+Although the urban tasks they explored differ from ours, their work can inspire ours to integrate LLMs to enhance fine-grained characterization capabilities.
+HREP (Zhou et al., 2023a) introduces prefix prompt to region embedding, but it lacks explicit extraction of task-adaptation.
+In this way, we propose both manually-designed and task-learnable prompts built on the graph-based pre-train model to address this issue.
+The categorization of related work is listed in Table 5.
+Table 5. Cats.of related works on urban region embedding.
+Category Non-Graph-based Graph-based
+Pretraining (Yan et al., 2024; Zhang et al., 2023; Li et al., 2023, 2024a; Xiao et al., 2024) (Fu et al., 2019; Zhang et al., 2021; Wu et al., 2022; Zhou et al., 2023a), GURPP (Ours)
+Prompting (Yan et al., 2024; Zhou et al., 2023a) GURPP (Ours)
 7. Conclusion and Future Works
@@
 Extensive experiments demonstrate the effectiveness of our approach.
-Future work may explore: 1) Integrating textual data and LLMs to uncover more universal urban knowledge; 2) Developing arbitrary region representations for flexible urban modeling; 3) Assessing the transferability of urban representations across cities to enhance general applicability.
+Future work may explore: 1) Integrating textual data and LLMs to uncover more universal urban knowledge; 2) Developing arbitrary region representations for flexible urban modeling (Sun et al., 2025); 3) Assessing the transferability of urban representations across cities to enhance general applicability.
 Appendix A Appendix
-A.1. Appendix of urban region graph construction
-Here we provide more details for the urban region graph construction.
-Specifically, we incorporate brand information for POI entities, such as Starbucks, to enhance the representation of POI functionalities.
-Altogether, we define eight node types shared across regions and cities: region, POI, POI category, brand, road, road category, junction, and junction category, comprising the node type set $T_{V}$.
-We also define six edge types, forming the edge type set $T_{E}$.
+A.1. Urban Region Graph Construction
+Here we provide more details for the urban region graph construction, with the schema shown in Figure 10.
+We define eight node types shared across regions and cities: region, POI, POI category, brand, road, road category, junction, and junction category in total, comprising the node type set $T_{V}$, and six edge types to form the edge type set $T_{E}$.
 These include the NearBy relationship, which connects region entities based on their geographical adjacency; the Contains relationship, which is based on spatial containment; the BrandOf relationship, which links brands to POI entities; and the CateOf, JCateOf, and RCateOf relationships, which are based on the category information of the entities.
-A.2. Appendix of region imagery encoding
-We supplement more details for the urban imagery encoding.
-Specifically, we employ ResNet (He et al., 2016) to extract visual features for each region image, such that $\vec{a}_{k}=\text{ResNet}(a_{k})$ and derive the comprehensive visual representation for the region by averaging the features of multiple images within the region.
-Then, a linear projection layer is utilized to project the visual feature into the representation space of region attribute, such that
-$\vec{h}_{r}^{img}=\text{Linear}(\frac{1}{m}\sum_{k=1}^{m}\vec{a}_{k}).$
-The contrastive loss $\mathcal{L}_{img}$ is then used to align the visual representation with the subgraph-centric region representation, ensuring a cohesive integration of visual and attribute information.
+Figure 10. Schema of urban region graph.
+A.2. Region Imagery Encoding
+We provide details on urban imagery encoding.
+ResNet (He et al., 2016) is utilized to extract visual features for each region image, represented as $\vec{a}_{k}=\text{ResNet}(a_{k})$.
+A comprehensive visual representation for the region is obtained by averaging features from multiple images: $\vec{h}_{r}^{img}=\text{Linear}(\frac{1}{m}\sum_{k=1}^{m}\vec{a}_{k}).$ We then adopt the contrastive loss $\mathcal{L}_{img}$ to align the visual representation with the subgraph-centric embedding.
 $\mathcal{L}_{img}$ $=\sum_{r\in\mathcal{R}}\Big{[}-\log\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)$
-$+\log\left(\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)+\sum_{j=1}^{B-1}\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{j}}{\tau}\right)\right)\Big{]}$
-Here, $B$ represents batch size and $\tau$ is the temperature parameter.
-A.3. Appendix of human mobility encoding
-Here we give a detailed description for human mobility encoding.
-To learn the embeddings, we reconstruct the departure and arrival flow distribution.
+$+\log\left(\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{r}^{img}}{\tau}\right)+\sum_{j=1}^{B-1}\exp\left(\frac{\vec{h}_{r}\cdot\vec{h}_{j}}{\tau}\right)\right)\Big{]},$
+where, $B$ represents batch size and $\tau$ is the temperature parameter.
+A.3. Mobility Pattern Encoding
+We provide a detailed description for mobility pattern encoding.
+We first reconstruct the departure and arrival flow distribution.
 Let $m_{ij}$ represent the number of trips from region $r_{i}$ to region $r_{j}$.
-The source and destination urban flow distributions are computed as $P^{{src}}(r_{j}|r_{i})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{ik}}$ and $P^{{dst}}(r_{i}|r_{j})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{kj}}$, respectively.
-We then reconstruct the source and destination distributions $\hat{P}^{{src}}(r_{j}|r_{i})$ and $\hat{P}^{{dst}}(r_{i}|r_{j})$ using the encoded inflow and outflow features:
-$\hat{P}^{{src}}(r_{j}|r_{i})=\frac{\exp((\vec{h}_{i}^{{src}})^{\top}\vec{h}_{j}^{{dst}})}{\sum_{k}\exp((\vec{h}_{i}^{{src}})^{\top}\vec{h}_{k}^{{dst}})}$
-$\hat{P}^{{dst}}(r_{i}|r_{j})=\frac{\exp((\vec{h}_{j}^{{dst}})^{\top}\vec{h}_{i}^{{src}})}{\sum_{k}\exp((\vec{h}_{j}^{{dst}})^{\top}\vec{h}_{k}^{{src}})}$
+The source and destination urban flow distributions are computed as $P^{{sc}}(r_{j}|r_{i})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{ik}}$ and $P^{{dt}}(r_{i}|r_{j})=\frac{m_{ij}}{\sum_{k=1}^{N}m_{kj}}$, respectively.
+We then reconstruct the source and destination distributions $\hat{P}^{{sc}}(r_{j}|r_{i})$ and $\hat{P}^{{dt}}(r_{i}|r_{j})$ using the encoded inflow and outflow features:
+$\hat{P}^{{sc}}(r_{j}|r_{i})=\frac{\exp((\vec{h}_{i}^{{sc}})^{\top}\vec{h}_{j}^{{dt}})}{\sum_{k}\exp((\vec{h}_{i}^{{sc}})^{\top}\vec{h}_{k}^{{dt}})},\hat{P}^{{dt}}(r_{i}|r_{j})=\frac{\exp((\vec{h}_{j}^{{dt}})^{\top}\vec{h}_{i}^{{sc}})}{\sum_{k}\exp((\vec{h}_{j}^{{dt}})^{\top}\vec{h}_{k}^{{sc}})}.$
 Finally, we learn the inflow and outflow embeddings by optimizing the following loss function:
-$\mathcal{L}_{flow}=\sum_{(r_{i},r_{j})\in\mathcal{R}\times\mathcal{R}}$ $-P^{src}(r_{j}|r_{i})\log\hat{P}^{src}(r_{j}|r_{i})$
-$-P^{dst}(r_{i}|r_{j})\log\hat{P}^{dst}(r_{i}|r_{j}),$
-which helps in accurately learning the inflow and outflow embeddings by comparing the true distributions with the reconstructed distributions.
-A.4. Appendix of attribute graph kernel function
+$\mathcal{L}_{flow}=\sum_{(r_{i},r_{j})}-P^{sc}(r_{j}|r_{i})\log\hat{P}^{sc}(r_{j}|r_{i})-P^{dt}(r_{i}|r_{j})\log\hat{P}^{dt}(r_{i}|r_{j}),$
+which aids in learning inflow and outflow embeddings by comparing true and reconstructed distributions.
+A.4. Attribute Graph Kernel Function
 We introduce how attribute graph kernel function $K_{p}(\cdot,\cdot)$ works.
-We first define the graph kernel function $K(\cdot,\cdot)$.
-The graph kernel function is a $P$-step random walk kernel (Nikolentzos and Vazirgiannis, 2020) to measure the similarity between graphs, which counts the number of common walks in both graphs within $P$ steps with its calculation as follows.
-Given $G_{1}=(V_{1},E_{1}),~{}G_{2}=(V_{2},E_{2})$, where $V$ and $E$ are the set of nodes and edges, respectively, we initially define the direct product graph $G_{\times}$ following the framework used in existing research (Nikolentzos and Vazirgiannis, 2020; Feng et al., 2022):
-(1) $G_{\times}=G_{1}\times G_{2}=(V_{\times},E_{\times})$
-where $V_{\times}=\{(v_{1},v_{2}):v_{1}\in V_{1},v_{2}\in V_{2}\}$, $E_{\times}=\{\{(v_{1},v_{2}),(u_{1},u_{2})\}:\{v_{1},u_{1}\}\in E_{1}\land\{v_{2},u_{2}\}\in E_{2}\}$.
-Following the theorem (Vishwanathan et al., 2010) that a single random walk on the direct product graph $G_{\times}$ is equivalent to a random walk on the two graphs $G_{1},~{}G_{2}$ at the same time, the $P$-step random walk kernel is computed in the following way:
-(2) $K(G_{1},G_{2})=\sum_{p=0}^{P}K_{p}(G_{1},G_{2})=\sum_{p=0}^{P}\lambda_{p}\sum_{i,j=1}^{|V_{\times}|}[A_{\times}^{p}]_{ij}$
+We choose a $P$-step random walk kernel (Feng et al., 2022) as graph kernel function $K(\cdot,\cdot)$ to measure the similarity between graphs, which counts the number of common walks in both graphs within $P$ steps.
+Given $G_{1}=(V_{1},E_{1}),~{}G_{2}=(V_{2},E_{2})$, where $V$ and $E$ are the set of nodes and edges, we initially define the direct product graph $G_{\times}$ (Feng et al., 2022): $G_{\times}=G_{1}\times G_{2}=(V_{\times},E_{\times}),$ where $V_{\times}=\{(v_{1},v_{2}):v_{1}\in V_{1},v_{2}\in V_{2}\}$, $E_{\times}=\{\{(v_{1},v_{2}),(u_{1},u_{2})\}:\{v_{1},u_{1}\}\in E_{1}\land\{v_{2},u_{2}\}\in E_{2}\}$.
+Following the theorem (Vishwanathan et al., 2010) that a single random walk on the direct product graph $G_{\times}$ is equivalent to random walk on the two graphs $G_{1},~{}G_{2}$ at the same time, the $P$-step random walk kernel is computed by:
+$K(G_{1},G_{2})=\sum_{p=0}^{P}K_{p}(G_{1},G_{2})=\sum_{p=0}^{P}\lambda_{p}\sum_{i,j=1}^{|V_{\times}|}[A_{\times}^{p}]_{ij},$
 where $A_{\times}$ is the adjacency matrix of the direct product graph, $\lambda_{p}$ is the sequence of weights and the $(i,j)$ item of $A_{\times}^{p}$ represents the $P$-length common walks between the $i$-th and $j$-th nodes.
 Extending the above definition to the scenario of graphs with node attributes, we use $\mathbf{X}\in\mathbb{R}^{n\times d}$ to denote the attribute matrix of a graph with $n$ nodes and attribute dimension $d$.
-Specifically, the node attribute matrices of $G_{1},G_{2}$ are denoted by $\mathbf{X}_{1}\in\mathbb{R}^{n_{1}\times d},\mathbf{X}_{2}\in\mathbb{R}^{n_{2}\times d}$.
+The node attribute matrices of $G_{1},G_{2}$ are denoted by $\mathbf{X}_{1}\in\mathbb{R}^{n_{1}\times d},\mathbf{X}_{2}\in\mathbb{R}^{n_{2}\times d}$.
 Thus, for the graph $G_{\times}=G_{1}\times G_{2}$, the node attribute matrix is $\mathbf{S}=\mathbf{X}_{1}\mathbf{X}_{2}^{\top},\mathbf{S}\in\mathbb{R}^{n_{1}\times n_{2}}$, where the item $(i,j)$ encodes the similarity between the $i$-th node in $G_{1}$ and the $j$-th node in $G_{2}$.
 We flattened $\mathbf{S}$ to $\mathbf{s}\in\mathbb{R}^{n_{1}n_{2}}$ to obtain the formula for the random walk kernel:
-(3) $K_{p}(G_{1},G_{2})=\sum_{i,j=1}^{|V_{\times}|}\mathbf{s}_{i}\mathbf{s}_{j}[A_{\times}^{p}]_{ij}=\mathbf{s}^{\top}A^{p}_{\times}\mathbf{s}.$
-A.5. Appendix of dataset statistics and sources
-We collect the datasets of region division, POI, urban imagery, taxi trips, road network, check-in, crime and crash for two representative cities in the United States: New York City (NYC) and Chicago (CHI).
+$K_{p}(G_{1},G_{2})=\sum_{i,j=1}^{|V_{\times}|}\mathbf{s}_{i}\mathbf{s}_{j}[A_{\times}^{p}]_{ij}=\mathbf{s}^{\top}A^{p}_{\times}\mathbf{s}.$
+A.5. Experiment Details
+A.5.1. Dataset Statistics and Sources.
+Statistics and sources of the datasets are shown in Table 6.
 The region division is based on census tracts and the urban imagery are 256 x 256 pixels with a resolution of 4.7 meters.
-The detailed statistics and sources of the dataset is shown in Table 5.
+For carbon emissions data provided in 1x1 $km$ grid cells, we map them to our regional divisions.
+If a region covers more than one grid cell, the emissions from all relevant cells are summed.
+For the income and education data, each region is assigned a specific value representing the average income or number of people with an education for that region.
 Table 6. Dataset Statistics and Sources.
@@
 POI 143,445 112,989 OSM (OpenStreetMap, 2024), Safegraph (Safegraph, 2024)
-Imagery 7,160 - ArcGIS (Arcgis, 2024), UUKG (Liu et al., 2023c)
+Imagery 7,160 - ArcGIS (Arcgis, 2024; Liu et al., 2023d)
 Taxi trip 12,684,353 3,889,032 NYCOD (NYC OpenData, 2024), CHIDP (Chicago Data Portal, 2024)
@@
 Crime 35,335 274,733 NYCOD (NYC OpenData, 2024), CHIDP (Chicago Data Portal, 2024)
-Check-in 106,902 - Foursquare (Yang et al., 2014; Foursquare, 2024)
+Check-in 106,902 - Foursquare (Foursquare, 2024)
 Crash - 150,535 CHIDP (Chicago Data Portal, 2024)
-A.6. Appendix of baselines description
-We compare our model with representative baseline models, including graph embedding methods, knowledge graph embedding methods and state-of-the-art region embedding methods.
+Carbon 1x1 $km$ 1x1 $km$ ODIAC (Oda and Maksyutov, 2015)
+Income 180 - NYCOD (NYC OpenData, 2024)
+Education - 869 CHIDP (Chicago Data Portal, 2024)
+A.5.2. Baselines Descriptions.
+Here we provide more details about the three types of baseline methods.
 I. Knowledge graph embedding methods:
-  • TransR (Lin et al., 2015) captures the complex interactions between entities and relationships in the knowledge graph by projecting them into different spaces and learning their representation in these spaces.
-  • TransR-N method uses the embedding of the region node learned by TransR as the region representation.
-  • TransR-G method is to average the embedding of the region subgraph learned by TransR as the region representation.
+  • $\circ$
+    TransR (Lin et al., 2015) projects entities and relationships into relation-specific spaces.
+TransR-N uses the region node embedding as the representation.
+TransR-G averages embeddings of all nodes in the region subgraph.
 II. Graph embedding methods:
-  • node2vec (Grover and Leskovec, 2016) generates a sequence of nodes by performing random walks on the graph to capture the neighborhood information of nodes, and then learns the representation of nodes using the Skip-Gram model.
-  • GAE (Kipf and Welling, 2016) consists of an encoder and a decoder.
-The encoder is used to learn the representation of nodes, and the decoder tries to reconstruct the adjacency matrix from these node representation.
+  • $\circ$
+    node2vec (Grover and Leskovec, 2016) uses random walks to generate node sequences and learns representations with the Skip-Gram model.
+  • $\circ$
+    GAE (Kipf and Welling, 2016) has an encoder for node representations and a decoder for reconstructing the adjacency matrix.
 III. State-of-the-art urban region embedding methods:
-  • MVURE (Zhang et al., 2021) constructs intra- and inter-region data multi-view based on human mobility and region attributes, and then performs multi-view fusion to learn region representation.
-  • MGFN (Wu et al., 2022) builds human mobility patterns and then learns intra- and inter-pattern relationships to obtain region representation through a multi-level cross-attention mechanism.
-  • HREP (Zhou et al., 2023a) designs a multi-relational heterogeneous graph of regions and used a relation-aware GCN to learn region representation, as well as a prefix prompt mechanism for downstream tasks.
-  • HRE is a part of HREP without prompt learning.
-  • ReCP (Li et al., 2024) is a multi-view comparative prediction model that learns region representation by capturing the unique information of a single view as well as the consistency between two views.
-A.7. Appendix of Implementation
-In our experiments, all the region embedding dimensions is set to 144.
-The number of HGT layer and head is 2 and 4, and the triplet margin parameter $\delta=2$, fusion loss weight $\mu=0.01$.
-We set the learning rate to 0.001 and the weight decay to 1e-6.
-In the task-learnable prompt module, we set the number of random walk kernel to 2 and the size of prompt graph to 6 and 8 of each layer.
-We implemented our model with Pytorch 1.8, and conducted experiments on a machine equipped with a NVIDIA RTX 3090 GPU, a 64 GB RAM and a 3.50 GHz CPU.
-A.8. Appendix of prompt graphs
-Figure 10 displays additional learned prompt graphs, with the majority of them being circular structural features.
-Figure 10. Visualization of additional learned prompt graphs.
+  • $\circ$
+    MVURE (Zhang et al., 2021) constructs multi-view data based on human mobility and region attributes for region representation.
+  • $\circ$
+    MGFN (Wu et al., 2022) learns relationships among human mobility patterns through cross-attention.
+  • $\circ$
+    HREP (Zhou et al., 2023a) creates a multi-relational graph and a relation-aware GCN with a non-graph prefix prompt. HRE is a variant without the prompt module.
+  • $\circ$
+    ReCP (Li et al., 2024a) learns region representation through a multi-view fusion that captures unique and consistent information.
+A.5.3. Implementation.
+We use five-fold cross-validation to evaluate GURPP and baselines on multiple tasks across both cities.
+In GURPP, the region embedding dimension is set to 144 (Zhou et al., 2023a).
+The HGT layer and head numbers are 2 and 4, respectively; triplet margin $\delta=2$; fusion loss weight $\mu=0.01$; learning rate is 0.001; and weight decay is 1e-6.
+In the task-learnable prompt module, we set the random walk kernel layers to 2, prompt graph sizes to 6 and 8, and region subgraph node size to 50.
+Due to the lack of satellite image datasets for CHI, we removed the image encoding module and adjusted the fusion layer dimensions accordingly.
+Our model is implemented with PyTorch 1.8 and DGL 2.2.1+cu118 on a machine with an NVIDIA RTX 3090 GPU, 64 GB RAM, and a 3.50 GHz CPU.
+Scalability experiments use an NVIDIA H100 80GB GPU, 1 TB RAM, and a 3.2 GHz CPU.
+A.5.4. Comparison with Language Prompt.
+We conduct a comparison with LLMs, including Deepseek R1, Gemini 2 Flash, in predicting the top 10 regions with the highest crime probabilities in NYC.
+The specific variant of GURPP used in this comparison is $\text{GURPP}_{M}$-P4.
+We provide each language model with detailed descriptions of the 180 regions in NYC, including region IDs and information about POIs.
+The prompt is as follows.
+I will provide information on the division of 180 regions in Manhattan, New York, with each region including a region ID and detailed description.
+Please analyze the potential crime risk of each area based on their descriptions, rank them according to the likelihood of criminal incidents from highest to lowest, and list the top 10 highest-risk regions.
+The regional information is formatted as: {mhtr0: mhtr0 is located in Manhattan, encompassing 1318 POIs, with the main types being Restaurants and Other Eating Places, Clothing Stores …} …